### v1

# bibliography.bib

Abstract
Given a pre-monotone Lagrangian link, we obtain Hofer energy estimates for Hamiltonian diffeomorphisms preserving it. Such estimates depend on the braid type of the Hamiltonian diffeomorphism only, and the natural language to talk about this phenomenon is provided by a family of norms on braid groups for surfaces with boundary. This generalises the results obtained by the first author to higher genus surfaces with boundary.
1 Introduction
Let $\Sigma_{g,p}$ be a compact surface with genus $g\geq 0$ and $p\geq 1$ boundary components (or punctures: we will use the two notions interchangeably since the content of the present paper is not affected by the difference). If $\Sigma_{g,p}$ is oriented, which will be the case throughout this paper, it is symplectic: there exists an area form $\omega$. We renormalise $\omega$ in such a way that $\int_{\Sigma_{g,p}}\omega=1$. Given a smooth function $H\in\mathcal{C}_{0}^{\infty}(\mathbb{S}^{1}_{t}\times\Sigma_{g,p};\mathbb{R})$ compactly supported in the interior of the surface for all times $t$, one may define its Hamiltonian vector field $X_{H_{t}}$ by
$\omega(X_{H_{t}},\cdot)=-dH_{t}$
The function $H$ one uses to define the Hamiltonian vector field is itself called Hamiltonian. The set of time 1-flows of such vector fields is a group, bearing the name of group of compactly-supported Hamiltonian diffeomorphisms of $\Sigma_{g,p}$. We denote it by $\text{Ham}_{c}(\Sigma_{g,p})$. If $H$ is a Hamiltonian, we write $\phi^{1}_{H}$ for its time 1-map.
The group $\text{Ham}_{c}(\Sigma_{g,p})$ comes with a natural metric on it, defined by Helmut Hofer in [hof90], as follows. For any smooth Hamiltonian $H$ on $\Sigma_{g,p}$ we define its oscillation
[…]
Let $\varphi\in\text{Ham}_{c}(\Sigma_{g,p})$: its Hofer norm is defined then as
$\lVert\varphi\rVert:=\inf_{H|\varphi=\phi^{1}_{H}}\lVert H\rVert_{(1,\infty)}$
[…]
$\text{If }\varphi,\varphi^{\prime}\in\text{Ham}_{c}(\Sigma_{g,p}),\,\,\,d_{H}(\varphi,\varphi^{\prime}):=\lVert\varphi(\varphi^{\prime})^{-1}\rVert$
It is rather complicated to provide examples of Hamiltonian diffeomorphisms with large Hofer norms. In principle, one expects Hofer energy to measure how “large” sets are moved around the symplectic manifold, and how complex their motion is. The first author of the present article constructed in [M] a class of compactly supported Hamiltonian diffeomorphisms of the disc whose Hofer energy satisfies some simple estimates from below. The goal is now to extend these estimates to higher genus symplectic surfaces.
1.1 Braid group on surfaces with genus and boundary
Geometrically speaking, the braid group can be seen as the fundamental group of the configuration space of the surface $\Sigma_{g,p}$. Let us fix $n\geq 1$ distinct points on $\Sigma_{g,p}$, $P_{1},\dots,P_{n}$, and consider the cartesian product $\left(\Sigma_{g,p}\right)^{n}$. The permutation group on $n$ letters, $\mathfrak{S}_{n}$ , acts on $\left(\Sigma_{g,p}\right)^{n}$, let $\text{Sym}^{n}\left(\Sigma_{g,p}\right)$ be the quotient. Denote by $\Delta$ the image of the points in $\left(\Sigma_{g,p}\right)^{n}$ for which at least two coordinates coincide. We define $\mathcal{B}_{n,g,p}$ as the fundamental group of the $n$-th (unordered) configuration space of $\Sigma_{g,p}$, with base point $\{P_{1},...,P_{n}\}$:
$\mathcal{B}_{n,g,p}:=\pi_{1}\left(\text{Sym}^{n}\left(\Sigma_{g,p}\right)\setminus\Delta\right)$
If $g=0,p=1$ (i.e. $\Sigma_{g,p}$ is the disc) we may describe this group via a presentation containing $n-1$ generators which obey the so-called braid relations:
$\mathcal{B}_{n,0,1}=\left\langle\sigma_{1},\dots,\sigma_{n-1}\middle|\begin{aligned} \sigma_{i}\sigma_{j}&=\sigma_{j}\sigma_{i}:|i-j|>1\\ \sigma_{i}\sigma_{i+1}\sigma_{i}&=\sigma_{i+1}\sigma_{i}\sigma_{i+1}:1\leq i\leq n-2\end{aligned}\right\rangle$
In such a case we have a remarkable homomorphism $\mathcal{B}_{n,0,1}\rightarrow\mathbb{Z}$, the linking number, obtained by sending every $\sigma_{i}$ to 1: this map will be denoted “$\mathrm{lk}$” throughout the paper. Note that each $\sigma_{i}$ corresponds geometrically to swapping $P_{i}$ and $P_{i+1}$ by a half-twist.
If however $g\geq 1,p\geq 1$, the presentation gets substantially more complicated. We are going to report here the results we need from [bel04]. Let us fix $n$ distinct base points on $\Sigma_{g,p}$. There exist then four families of generators (for a picture, see Figure 1):
  • $\sigma_{1},\dots,\sigma_{n-1}$: they correspond to the generators of $\mathcal{B}_{n,0,1}$, i.e. to half-twist swapping two of the base points, in such a way that the images of the paths are contained in a disc on the surface;
  • $a_{1},\dots,a_{g}$, $b_{1},\dots,b_{g}$: they are obtained by homologically independent loops based at the first base point, $P_{1}$ (the other base points are instead fixed throughout the path). One way to describe them is seeing $\Sigma_{g,p}$ as the connected sum of $g$ tori with $p$ punctures, so that the loops $a_{i}$ and $b_{i}$ represent the generators of the homology of the $i$-th torus;
  • $z_{1},\dots,z_{p-1}$: they correspond to loops based at $P_{1}$ and winding around the $i-th$ puncture exactly once (here as well the other base points are fixed). Remark that there are $p-1$ generators, but $p$ punctures. A loop around the last puncture may in fact be written as composition of the others.
The relations in this group are rather complex, and we are not going to report them here to keep the presentation lean. The interested reader will find a detailed account in [bel04, Theorem 1.1].
1.2 The main result
Let $\underline{L}=L_{1},\dots,L_{k+g}$ be a set of $k+g$ disjoint, embedded circles on $\Sigma_{g,p}$: we call $\underline{L}$ a Lagrangian link.
Definition 1.1.
We say that $\underline{L}$ is premonotone if the following conditions are satisfied:
  • i)
    exactly $g$ of the circles in $\underline{L}$ are non contractible;
  • ii)
    $\Sigma_{g,p}\setminus\bigcup_{i=1}^{k+g}L_{i}$ is a disjoint union of $k$ discs $(B_{j})_{j=1,\dots,k}$ and a pair of pants $B_{k+1}$ with $p+k+2g-1$ legs;
  • iii)
    There exist $\lambda\in\left(\frac{1}{k+1},\frac{1}{k}\right)$, such that $\lambda=\int_{B_{j}}\omega$ for $j=1,\dots,k$.
Given a premonotone Lagrangian link, we denote by $\text{Ham}_{\underline{L}}(\Sigma_{g,p})$ the subgroup of $\text{Ham}_{c}(\Sigma_{g,p})$ stabilising the Lagrangian link as a set:
$\text{Ham}_{\underline{L}}(\Sigma_{g,p}):=\Set{\varphi\in\text{Ham}_{c}(\Sigma_{g,p})}{\exists\sigma\in\mathfrak{S}_{k+g},\varphi(L_{i})=L_{\sigma(i)}}$
To a diffeomorphism in $\text{Ham}_{\underline{L}}(\Sigma_{g,p})$ we may associate an element of the braid group $\mathcal{B}_{k,g,p}$. This function is defined the following way: choose one base point per contractible circle in $\underline{L}$, denote it by $P_{i}\in L_{i}$ for $i=1,\dots,k$. Any Hamiltonian isotopy $(\varphi_{t})_{t\in[0,1]}$ between the identity and $\varphi\in\text{Ham}_{\underline{L}}(\Sigma_{g,p})$ provides then a collection of $k$ curves $t\mapsto\varphi_{t}(P_{i})$. Remark that there exists a permutation $\sigma$ in $\mathfrak{S}_{k+g}$ such that for all $i$, $\varphi(L_{i})=L_{\sigma(i)}$, and contractible circles are mapped to contractible circles. For each contractible $L_{i}$, choose then a path in $L_{\sigma(i)}$ connecting $\varphi(P_{i})$ to $P_{\sigma(i)}$. The concatenation of the curves $t\mapsto\varphi_{t}(P_{i})$ with the respective connecting paths yields an element in $\mathcal{B}_{k,g,p}$. The braid obtained this way does not depend on the choice of the base points, of connecting paths or of the Hamiltonian isotopy (a proof of this fact is included later). We thus have defined a groups homomorphism
$b:\text{Ham}_{\underline{L}}(\Sigma_{g,p})\rightarrow\mathcal{B}_{k,g,p}$
Denote by $\mathcal{B}_{\underline{L}}$ the image of $b$, which will be described in detail in Section 3.2.
Remark 1.1.
If $g\geq 1$, the braid type function $b$ cannot be surjective. Indeed, any non contractible circle in $\underline{L}$ is Hamiltonianly non-displaceable, and therefore has to be mapped to itself by any element in $\text{Ham}_{\underline{L}}(\Sigma_{g,p})$. In the genus 0 case $b$ is instead surjective, see [M].
Our main result is the following:
Theorem 1.2.
There exists a family $(\mathfrak{f}_{(v_{1},v_{2})})_{v_{i}\in V}$ of group homomorphisms $\mathcal{B}_{\underline{L}}\rightarrow\mathbb{R}$ indexed on the square of a $p$-dimensional simplex $V$ (defined in Section 4.2) such that for every $v_{i}\in V$ and $\varphi,\psi\in\text{Ham}_{\underline{L}}(\Sigma_{g,p})$,
$d_{H}(\varphi,\psi)\geq\frac{1}{2}\lvert\mathfrak{f}_{(v_{1},v_{2})}(b(\varphi\psi^{-1}))\rvert$ (1)
In particular, we have the following estimate for the Hofer norm:
$\lVert\varphi\rVert\geq\frac{1}{2}\sup\limits_{(v_{1},v_{2})\in V^{2}}\lvert\mathfrak{f}_{(v_{1},v_{2})}(b(\varphi))\rvert$ (2)
Let us define the function on the image of $b$
$\mathfrak{f}(g)=\max_{(v_{1},v_{2})\in V^{2}}\lvert\mathfrak{f}_{(v_{1},v_{2})}(g)\rvert$ (3)
In the case of the generators $\sigma_{j}$, $a_{i}$, $b_{i}^{-1}a_{i}b_{i}$ we then obtain the explicit values:
$\mathfrak{f}(\sigma_{j})=\frac{1}{2}\mathfrak{f}(a_{i})=\frac{1}{2}\mathfrak{f}(b_{i}^{-1}a_{i}b_{i})=\frac{1}{2(k+g)}\frac{(k+1)\lambda-1}{k+2g-1}$
Remark 1.3.
We compute in (24) explicit values for the maximum in 3
Remark 1.4.
This result seems counter-intuitive when compared to a result of Khanevsky’s contained in [kha15]. In that paper, Khanevsky defines the notion of homological trajectory of a Hamiltonian diffeomorphism defined on a surface and fixing a given disc. He then proceeds to prove that, whenever the genus of the surface is positive, there exists a constant only depending on the disc such that one can realise any trajectory with a Hamiltonian diffeomorphism of Hofer energy lower than this threshold. In fact, it is easily seen that such diffeomorphisms cannot stabilise any non-contractible component of the Lagrangian link $\underline{L}$, therefore they do not appear in our treatment. If anything, this shows that to find a result like ours, the constraint of fixing a number of non-contractible circles equal to the genus is in fact minimal (otherwise it is possible to construct those diffeomorphisms, realising arbitrarily complex braids with bounded Hofer energy).
We proceed now to define a norm on $\mathcal{B}_{\underline{L}}$. For any $g\in\mathcal{B}_{\underline{L}}$ we set
$\lVert g\rVert_{\underline{L}}:=\inf_{\varphi\in\text{Ham}_{\underline{L}}(\Sigma_{g,p}),b(\varphi)=g}\lVert\varphi\rVert$ (4)
[…]
A corollary of the main result from this paper is an estimate from below of $\lVert\cdot\rVert_{\underline{L}}$:
Corollary 1.5.
Let $g\in\mathcal{B}_{\underline{L}}$. Then, if $\mathfrak{f}$ is the function from the main result,
$\lVert g\rVert_{\underline{L}}\geq\frac{1}{2}\cdot\lvert\mathfrak{f}(g)\rvert$ (5)
As in [M], this is not enough to conclude non-degeneracy as soon as $k+g\geq 2$ (as soon as the genus is positive, there are no interesting cases we can consider for $k+g\leq 2$). We may however adapt a proof by Chen contained in [chen23] to show that
Theorem 1.6.
For a pre-monotone $\underline{L}$, the pseudonorm $\lVert\cdot\rVert_{\underline{L}}$ is non-degenerate.
Chen had in particular proved this fact for braids on the disc.
Recall the existence of the real-valued Calabi morphism on $\text{Ham}_{c}(\Sigma_{g,p})$ (for a definition, see for instance [mcDSal16, Section 10.3]). An immediate corollary of this result is the existence of Hamiltonian diffeomorphisms in the kernel of Calabi with non-zero asymptotic Hofer norm.
For any $\varphi\in\text{Ham}_{c}(\Sigma_{g,p})$, we define its asymptotic Hofer norm by
$\lVert\varphi\rVert_{\infty}:=\lim_{n}\frac{\lVert\varphi^{n}\rVert}{n}$
We may now state the following elementary consequence of Theorem 1.2:
Corollary 1.7.
Let $\varphi\in\text{Ham}_{\underline{L}}(\Sigma_{g,p})$. Then
$\lVert\varphi\rVert_{\infty}\geq\frac{1}{2}\cdot\lvert\mathfrak{f}(b(\varphi))\rvert$
Moreover, any braid type in the image of $b$ may be realised by a diffeomorphism in the kernel of Calabi.
Proof.
Since $\mathfrak{f}_{(v_{1},v_{2})}$ and $b$ are both morphisms of groups for all $v_{i}\in V$, the only statement which is not entirely trivial is the one about the kernel of Calabi. This is easily seen, since Calabi is a morphism of groups, and we may compose any element in $\text{Ham}_{\underline{L}}(\Sigma_{g,p})$ by a diffeomorphism realising the trivial braid and arbitrary Calabi. ∎
1.3 Structure of the paper
We are going to start our paper by recalling the necessary notions from the construction of Quantitative Heegaard-Floer homology as in [CGHMSS22]. We end Section 2 by recalling the Künneth isomorphism in [MT] which will be used in a simplified proof of a particular case of our result (this may be found at the end of the paper and is independent of the rest). In this section we are also going to prove monotonicity of Lagrangian links under more general hypotheses than those in [CGHMSS22] (we need this fact to be able to generalise the methods of [M] to surfaces with genus). After reviewing this construction, we give some necessary details about braid and Hamiltonian diffeomorphism groups, details which are used in the construction of the Hofer-Lipschitz functions and in the “braid type” function
$b:\text{Ham}_{\underline{L}}(\Sigma_{g,p})\rightarrow\mathcal{B}_{k,g,p}$
It is here that we give a characterisation of the image of $b$ using the description in [bel04].
After this necessary introduction, we go on to construct the family of the Hofer Lipschitz functions $\mathfrak{f}_{(v_{1},v_{2})}$, and explain the optimisation process which yields the function $\mathfrak{f}$ from the main result. In this section we also describe in detail the intersection phenomena taking place in the symmetric product. We end by showing how the proofs contained in [chen23] also apply in our setting, and prove non-degeneracy of the norms on the braid groups of surfaces we defined above. The paper closes with a simplified proof of our result in which we only consider Hamiltonian diffeomorphisms with compact support in a fixed disc. This section will only rely on the ideas from [M] and the Künneth isomorphism in [MT].
Acknowledgments
The first author would like to thank Vincent Humilière for his support and interest in the project, Pierre Godfard for office chats about symmetric products of Riemann surfaces, Paolo Bellingeri and Cheuk Yu Mak for their helpful indications. Fabio Gironella’s patient Inkscape tutorial also saved him a lot of time. The second author would also like to thank Vincent Humilière, Cheuk Yu Mak as well as his advisor Sobhan Seyfaddini for his support and helpful suggestions.
2 Quantitative Heegaard-Floer homology and a Künneth Formula
We recall the construction of Heegaard-Floer homology and link spectral invariants. We are mainly following [CGHMSS22, Section 6], although we are considering homology instead of cohomology, and our conventions differ slightly from theirs. Moreover, we weaken the necessary conditions for the definition of the homology in the case of surfaces with genus.
Let $\Sigma_{g}$ be a closed, symplectic surface, and $\underline{L}$ be a collection of $k+g$ circles satisfying the conditions detailed in Definition 1.1 setting $p=0$. Let $\varphi\in\text{Ham}(\Sigma_{g})$ be a Hamiltonian diffeomorphism defined by a Hamiltonian $H$, and such that $\forall i,j=1,\dots,k+g$, $\varphi(L_{i})\pitchfork L_{j}$.
The symmetric group on $k+g$ letters acts on $(\Sigma_{g})^{k+g}$ by permutation of the coordinates. We denote indifferently by $\text{Sym}^{k+g}(\Sigma_{g})$ or $X$ the quotient of the cartesian product by this action: this set may be given the structure of a symplectic manifold, as in [Perutz]. We consider then the image of the Lagrangian torus $\prod_{i=1}^{k+g}L_{i}$ under the quotient map: since $\prod_{i=1}^{k+g}L_{i}$ does not intersect the diagonal (the subset of the cartesian product where two or more coordinates coincide), the quotient map induces a diffeomorphism between $\prod_{i=1}^{k+g}L_{i}$ and its image, which we denote by $\text{Sym}^{k+g}(\underline{L})$. Let also $\Delta$ be the image of the diagonal under the quotient map in $\text{Sym}^{k+g}\Sigma_{g}$. The push-forward of the product symplectic form on $\Sigma_{g}^{k+g}$, which we denote by $\omega_{X}$, is not smooth near $\Delta$: in [Perutz] the author applies a convolution to smoothen it. The smoothened form coincides with $\omega_{X}$ outside an arbitrarily small neighbourhood of the diagonal, so that $\text{Sym}^{k+g}(\Sigma_{g})$ turns out to be Lagrangian. We shall still denote the smoothened form by $\omega_{X}$: this abuse of notation will not affect any of the following.
In a similar way, to $\varphi$ we may associate a Hamiltonian diffeomorphism of $\text{Sym}^{k+g}\Sigma_{g}$: it is going to be defined by approximating the function
$\text{Sym}H_{t}:\text{Sym}^{k+g}\Sigma_{g}\rightarrow\mathbb{R},\,\,\,\text{Sym}H_{t}([x_{1},\dots,x_{k+g}]):=\sum_{i=1}^{k+g}H_{t}(x_{i})$
via another one which is a time-dependent constant near $\Delta$, and coincides with $H_{t}$ away from it. By an abuse of notation we denote by $\text{Sym}^{k+g}(\varphi)$ the diffeomorphism generated by an approximation of the quotient Hamiltonian (none of the following constructions will depend on the actual approximation we choose). The Quantitative Heegaard-Floer Homology of the pair $(\underline{L},H)$ is defined to be the traditional Lagrangian Floer Homology in $\text{Sym}^{k+g}\Sigma_{g}$ of the pair $(\text{Sym}^{k+g}(\underline{L}),\text{Sym}^{k+g}(\varphi))$. The homology is well defined using Proposition 2.1 and the arguments contained in [CGHMSS22]: our result is needed since in [CGHMSS22] they need more non contractible link components to build the theory. We may define the Floer complex
$CF(\underline{L},H)$
To define the complex we consider a point $x\in\text{Sym}^{k+g}(\underline{L})$ as a constant path. Generators of the Floer complex are Hamiltonian chords from $\text{Sym}^{k+g}(\underline{L})$ to $\text{Sym}^{k+g}(\varphi)\text{Sym}^{k+g}(\underline{L})$ which are homotopic to the reference path $x$ A capping is the choice of a homotopy between the Hamiltonian chord and $x$. The differential of the complex is defined by counting appropriate pseudo-holomorphic discs in $X$ with boundary conditions on the Lagrangian tori $\text{Sym}^{k+g}(\underline{L})$ and $\text{Sym}^{k+g}(\varphi)\text{Sym}^{k+g}(\underline{L})$. For details, we refer to [CGHMSS22] (our conventions here are consistent with those of [MT]).
The complex is filtered by the symplectic action: given a Hamiltonian chord $y$ with capping $\hat{y}$, its action is
[…]
The very last term is the intersection product between the capping and the diagonal: this is well defined because the capping sends the boundary of the square to the Lagrangians, which do not intersect the diagonal by assumption. The group $\pi_{2}(X)$ acts on the space of capped intersections, given by recapping: since the Lagrangian $\text{Sym}^{k+g}(\underline{L})$ is monotone, the symplectic action is shifted by multiples of $\lambda$ by recapping.
The homology of $CF(\underline{L},H)$ is non-zero (this follows from an application of the Künneth formula together with continuation maps for the homology, see Section 2.3) and admits a product structure with unit $\mathbf{1}$: we define the spectral invariant $c_{\underline{L}}(H)$ as:
$c_{\underline{L}}(H):=\frac{1}{k+g}\inf\Set{\alpha\in\mathbb{R}}{\mathbf{1}\text{ may be represented by a chain of action}<\alpha}$
If the Hamiltonian $H$ is mean-normalised in fact $c_{\underline{L}}(H)$ is an invariant of the homotopy class of the Hamiltonian path $t\mapsto\varphi_{t}$ with fixed endpoints. If $g\geq 1$, the group of Hamiltonian diffeomorphisms is simply connected, so that $c_{\underline{L}}$ provides an invariant of the time 1-map. The resulting map
$c_{\underline{L}}:\text{Ham}(\Sigma_{g})\rightarrow\mathbb{R}$
is moreover Hofer Lipschitz. The reason why $c_{\underline{L}}$ is defined on the whole of $\text{Ham}(\Sigma_{g})$ and not just on the non-degenerate diffeomorphisms is given by Hofer-density of the latter set, and by the Hofer Lipshitz property.
2.1 Sketch of the proof when the genus is 0
We are going to recall the main ideas behind the proof in [M].
In the class of Hamiltonian diffeomorphisms $\text{Ham}_{\underline{L}}(\Sigma_{g,p})$, the reference path we use for the definition of $CF(\underline{L},H)$ is essentially a representative of the braid class of $b(\varphi)$. If $\hat{y}$ is a capped intersection point, the intersection product $[\hat{y}]\cdot\Delta$ can be connected to the linking number of $b(\varphi)$: we aim to then isolate this contribution in the expression of the action.
We are lead to consider a difference of spectral invariants: we embed the disc $\mathbb{D}$ symplectically into spheres of areas $1+s_{i}$ ($i=1,2$), with $10\leq s_{i}\leq(k+1)A-$ 1, in such a way that the obtained Lagrangian links are monotone with same $\lambda$ but different $\eta$. We thus want to compute
$c_{\underline{L}_{s_{1}}}(\varphi)-c_{\underline{L}_{s_{2}}}(\varphi)$ (7)
where the $\underline{L}_{s_{i}}$ are the two obtained Lagrangian links.
We define a biholomorphism between the symmetric products: this induces a chain isomorphism between the two $CF(\underline{L}_{s_{i}},H)$, but such isomorphism does not respect the action filtration. The difference (7) is exactly the failure to respect the action filtration of this chain isomorphism. This means that the spectral invariants in (7) may be computed using capped intersection points related by the biholomorphism. Since the two Lagrangian links share furthermore the same parameter $\lambda$, we may change the capping of this intersection point freely; without loss of generality therefore we may assume that such capping is contained in $\text{Sym}^{k}\mathbb{D}\subset\text{Sym}^{k}\mathbb{S}^{2}$. The proof ends showing that under this assumption
$c_{\underline{L}_{s_{1}}}(\varphi)-c_{\underline{L}_{s_{2}}}(\varphi)=(\eta_{s_{2}}-\eta_{s_{1}})[\hat{y}]\cdot\Delta$
showing that the intersection product coincides with the linking number up to a factor of $-2$. Choosing optimal $s_{i}$ and applying the Hofer Lipschitz property of the difference in (7) yields the main result of [M].
2.2 Proof of monotonicity
To show that the differential on the complex is well defined, we need to be able to apply the Gromov compactness to the moduli spaces. This is done in [CGHMSS22] by proving a monotonicity property which gives a uniform bound to the energy of Maslov index 1 strips with given asymptotics. However, they only prove the property for links such that the closure of each connected component of the complement is planar. In this section, we prove the monotonicity property for a larger class of links, which in turn imply that the Floer complex is well defined.
For more generality, we assume that $\int_{\Sigma_{g}}\omega=A$ (the area $A$ is not necessarily 1). In the following, we split the circles in $\underline{L}$ in two classes: the contractible and the non-contractible ones. We assume that $L_{1},...,L_{k}$ ($k\geq 2$) bound disjoint discs $B_{1},...,B_{k}$ of area $\lambda\in[\frac{A}{k+1},\frac{A}{k})$, and that $\alpha_{1},...,\alpha_{g}$ ($\alpha_{i}:=L_{k+i}$) be meridians for each handle of $\Sigma_{g}$. This way, we have $\underline{L}=L_{1}\cup...\cup L_{k}\cup\alpha_{1}\cup...\cup\alpha_{g}$ and $X:=\text{Sym}^{k+g}(\Sigma_{g})$. Let $B_{k+1}$ be the only connected component of $\Sigma_{g}\setminus\underline{L}$ that is not a disc, and let $A_{k+1}$ be its area. Let $\eta$ be the real number satisfying
$A_{k+1}+2\eta(k+2g-1)=\lambda$
Then $\eta$ can be recovered by the formula
$\eta=\frac{\lambda-A_{k+1}}{2(k+2g-1)}=\frac{(k+1)\lambda-A}{2(k+2g-1)}$
and in particular is non-negative.
Proposition 2.1.
For all $[u]$ in the image of $\pi_{2}(X,\text{Sym}^{k+g}(\underline{L}))\to H_{2}(X,\text{Sym}^{k+g}(\underline{L}))$ one has
$\omega_{X}([u])+\eta\Delta\cdot[u]=\frac{\lambda}{2}\mu([u])$
Proof.
For $11\leq i\leq k+$ 1, let $\overline{B}_{i}$ be the closure of $B_{i}$ in $\Sigma_{g}$. Let $k_{i}$ be the number of boundary components of $\overline{B}_{i}$. For each $i$, fix a point $a_{i}$ in $B_{i}$, and let $X_{a_{i}}$ be the projection of $\Sigma_{g}^{k+g-1}\times\{a_{i}\}$ in $\text{Sym}^{k+g}(\Sigma_{g})$. In [CGHMSS22, Section 4.5], it is explained how, when $\overline{B}_{i}$ is planar, one can construct a disc class $[u_{i}]$ in $H_{2}(X,\text{Sym}\underline{L})$, which satisfies $[u_{i}]\cdot X_{a_{j}}=\delta_{i,j}$, $[u_{i}]\cdot\Delta=k_{i}-1$, and $\mu([u_{i}])=2$. Here, the $\overline{B}_{i}$’s for $1\leq i\leq k$ are discs, therefore we can apply this construction to get $k$ classes satisfying $\omega_{X}([u_{i}])+\eta\Delta\cdot[u_{i}]=\lambda+0=\frac{\lambda}{2}\mu([u_{i}])$.
However, $\overline{B}_{k+1}$ is not a planar domain (it is equal to the whole surface minus $k$ disjoint discs). But we can still apply a similar construction:
Let $\widehat{D}:=\overline{B}_{k+1}\sqcup\coprod_{1\leq j\leq g}D_{j}$, where the $D_{j}$’s are copies of the closed unit disc D in $\mathbb{C}$. Let $\pi_{\widehat{D}}:\widehat{D}\to D$ be a $(k+g)$-fold simple branched covering, such that $\pi_{\widehat{D}}|_{D_{j}}$ is a biholomorphism for all $j$, and $\pi_{\widehat{D}_{i}}|_{\overline{B}_{k+1}}$ is a topological $k$-fold simple branched cover, such that $a_{k+1}$ is not a branched point. Since $k\geq 2$, such a branched cover always exists; one can construct it using techniques presented in John Etnyre’s lecture notes [Etnyre]. Let $v_{k+1}:\widehat{D}\to\Sigma_{g}$ be a map whose restriction to $\overline{B}_{k+1}$ is the identity, and which sends $D_{j}$ to a point in $\alpha_{j}$. Then, by tautological correspondence, we get a map $u_{k+1}:(D,\partial D)\to(X,\text{Sym}\underline{L})$, defined by $u_{k+1}(z)=v_{k+1}(\pi_{\widehat{D}}^{-1}(z))$. The class $[u_{k+1}]$ also satisfies $[u_{k+1}]\cdot X_{a_{i}}=v_{k+1}\cdot a_{i}=\delta_{i,k+1}$ for $11\leq i\leq k+$ 1. Therefore, the proof of [CGHMSS22, Lemma 4.10] goes through, and we get that the image of $\pi_{2}(X,\text{Sym}\underline{L})\to H_{2}(X,\text{Sym}\underline{L})$ is freely generated by $\{[u_{i}]\}_{i=1}^{k+1}$, and that the image of $\pi_{2}(X)\to H_{2}(X,\text{Sym}\underline{L})$ is freely generated by $\sum_{i=1}^{k+1}[u_{i}]$.
Since the $[u_{i}]$ satisfy the equation $\omega_{X}([u_{i}])+\eta\Delta\cdot[u_{i}]=\frac{\lambda}{2}\mu([u_{i}])$ for $1\leq i\leq k$, it only remains to show it for $[u_{k+1}]$.
We start by computing $[u_{k+1}]\cdot\Delta$ in a similar fashion as in [CGHMSS22]. Let $\pi:\Sigma_{g}\to S^{2}$ be a topological $k+g$-fold simple branched covering of the sphere. Let $u:S^{2}\to X$ be the map tautologically corresponding to the pair $(\pi,id_{\Sigma_{g}})$. Then, since $[u]$ is in the image of $\pi_{2}(X)\to H_{2}(X,\text{Sym}\underline{L})$, there exists an integer $c$ such that $[u]=c\sum_{i=1}^{k+1}[u_{i}]$. Since $[u]\cdot X_{a_{i}}=1$ for all $i$, we get that $c=1$. Therefore, $[u]\cdot\Delta=(\sum_{i=1}^{k+1}[u_{i}])\cdot\Delta=[u_{k+1}]\cdot\Delta$. Since $id_{\Sigma_{g}}$ is injective, there is a one-to-one correspondence between branched point of $\pi$ and points of $S^{2}$ whose image by $u$ lies in $\Delta$. Moreover, since the branched points are simple, we have that $[u]\cdot\Delta$ is actually equal to the number of branched points of $\pi$. Therefore, the Riemann-Hurwitz formula implies that $2-2g=2(k+g)-[u]\cdot\Delta$, and we finally get $[u_{k+1}]\cdot\Delta=2(k+2g-1)$. It only remains to compute the Maslov index of $[u_{k+1}]$:
$\mu([u_{k+1}])=\mu\left([u]-\sum\limits_{i=1}^{k}[u_{i}]\right)=2\langle c_{1}(TX),[u]\rangle-2k$
Let $v$ be the Abel-Jacobi map, from $\text{Sym}^{d}(\Sigma_{g})$ to its Jacobian variety $J$, which is isomorphic to the $2g$-dimensional torus $T^{2g}$. According to [Arbarello, Chapter VII, Section 5], there is a class $\theta$ in $H^{2}(J)$ and a point $q$ in $\Sigma_{g}$ such that:
$c_{1}(T\text{Sym}^{d}\Sigma_{g})=(d-g+1)PD(X_{q})-v^{*}\theta$
Now, in our case, $d=k+g$, and as above $[u]$ is a generator of the image of $\pi_{2}(X)\to H_{2}(X,\text{Sym}\underline{L})$. Since $J=T^{2g}$ is aspherical, $v_{*}[u]$ vanishes, and
$\mu([u_{k+1}])=2\langle c_{1}(TX),[u]\rangle-2k=2(k+1)[u]\cdot X_{q}-2k=2$
where the last equality comes from the fact that $[u]\cdot X_{q}=id_{\Sigma_{g}}\cdot q=1$ when $q$ is not a branched point of $\pi$, which we can assume by perturbing $\pi$ if necessary. As claimed, we get
$\omega_{X}([u_{k+1}])+\eta\Delta\cdot[u_{k+1}]=A_{k+1}+2\eta(k+2g-1)=\lambda=\frac{\lambda}{2}\mu([u_{i}])$
Remark 2.2.
One can use the results of this section to show the monotonicity property for a more general class of links. In fact, let $\underline{L}$ be a link on $\Sigma_{g}$, and $B_{1},...,B_{s}$ be the connected components of $\Sigma_{g}\setminus\underline{L}$. Let $k_{i}$ be the number of boundary components of $\overline{B}_{i}$ (the closure of $B_{i}$ in $\Sigma_{g}$), $g_{i}$ be its genus, and $A_{i}$ be its area. Assume the following is satisfied:
  • for all $i$, $\overline{B}_{i}$ contains exactly one non-contractible component of $\underline{L}$ around each of its handles (otherwise said, the non-contractible components of the link generate a $g$-dimensional subspace of the first homology of $\Sigma_{g}$);
  • if $g_{i}>0$, $k_{i}\geq 2$;
  • there exists constants $\lambda$ and $\eta$ such that for all $1\leq i\leq s$, $A_{i}+2\eta(k_{i}+2g_{i}-1)=\lambda$.
Then, $\underline{L}$ satisfies the monotonicity property: for all $[u]$ in the image of
$\pi_{2}(X,\text{Sym}(\underline{L}))\to H_{2}(X,\text{Sym}(\underline{L}))$
one has $\omega_{X}([u])+\eta\Delta\cdot[u]=\frac{\lambda}{2}\mu([u])$. As in [CGHMSS22], it is enough to check the monotonicity property on the basic disc classes tautologically corresponding to the $\overline{B}_{i}$’s. When they are planar ($g_{i}=0$), it is shown in [CGHMSS22] that they satisfy the monotonicity property by embedding $\overline{B}_{i}$ into the sphere. If $g_{i}\neq 0$, we can embed $\overline{B}_{i}$ into $\Sigma_{g_{i}}$ by gluing $k_{i}$ discs of area $\lambda$ on the boundary components. Then, we can apply the result of this section to get the monotonicity property for $\overline{B}_{i}$.
2.3 The Künneth formula
In this section we present a result shown by Cheuk Yu Mak and the second author in [MT]. We will use it in Section 6 to give an easier proof of our main result in a simpler case. The setting is the following: let $\underline{L}$ and $\underline{K}$ be two transverse $\eta$-monotone Lagrangian links on a closed surface $(\Sigma,\omega)$. Let $(E,\omega_{E})$ be a two-dimensional torus, $\alpha$ be a non-contractible circle on $E$, and $\alpha^{\prime}$ be a small Hamiltonian deformation of $\alpha$, transverse to $\alpha$. Pick a point $\sigma_{1}\in\Sigma$ away from $\underline{L}$ and $\underline{K}$, and a point $\sigma_{2}\in E$ away from the support of the isotopy between $\alpha$ and $\alpha^{\prime}$. Pick open discs $B(\sigma_{1})$ and $B(\sigma_{2})$ that contain $\sigma_{1}$ and $\sigma_{2}$ respectively, and that are also away from the Lagrangian links and the isotopy. Let $\Sigma^{\prime}:=\Sigma\#E$ be the connected sum of $\Sigma$ and $E$, obtained by gluing together $\Sigma\setminus\{\sigma_{1}\}$ and $E\setminus\{\sigma_{2}\}$ along the two cylinders $B(\sigma_{1})\setminus\{\sigma_{1}\}$ and $B(\sigma_{2})\setminus\{\sigma_{2}\}$. We equip $\Sigma^{\prime}$ with an area form $\omega^{\prime}$ which coincides with $\omega$ on $\Sigma\setminus B(\sigma_{1})$, with $\omega_{E}$ on the support of the isotopy between $\alpha$ and $\alpha^{\prime}$, and such that $\omega^{\prime}(\Sigma^{\prime})+4\eta=\omega(\Sigma)$. (For this last condition to be satisfied, we need to assume that the support of the isotopy between $\alpha$ and $\alpha^{\prime}$ has small enough area.) The result is the following ([MT, Theorem 6]): for an appropriate choice of complex structure, there is a filtered chain complex isomorphism
$CF(\underline{L},\underline{K})\otimes CF(\alpha,\alpha^{\prime})\xrightarrow{\sim}CF(\underline{L}\cup\alpha,\underline{K}\cup\alpha^{\prime})$
In this Künneth formula for the connected sum, $\underline{L}$ and $\underline{K}$ are viewed as links on $\Sigma$, $\alpha$ and $\alpha^{\prime}$ are Lagrangians on $E$, and $\underline{L}\cup\alpha$ and $\underline{K}\cup\alpha^{\prime}$ are links in $\Sigma^{\prime}$. The proof of this statement is inspired by the “stabilisation invariance” shown in [OS, Section 10]. The first step is to identify the generators of those vector spaces. Then to show that we get a chain isomorphism, one needs to identify the moduli spaces counting curves contributing to the differentials on both sides. This is done by a gluing argument and Gromov compactness. Finally, the part of the statement concerning the action filtration is a straightforward consequence of our choice of symplectic form $\omega^{\prime}$.
Remark 2.3.
Compared to [MT] who showed the isomorphism in the case $\eta=0$, we needed here to add the condition $\omega^{\prime}(\Sigma^{\prime})+4\eta=\omega(\Sigma)$ to make sure that the action filtration is preserved. This equality implies that $\underline{L}\cup\alpha$ and $\underline{K}\cup\alpha^{\prime}$ are also $\eta$-monotone links on $(\Sigma^{\prime},\omega^{\prime})$.
We can use this Künneth isomorphism to show that the Heegaard Floer homology of a monotone link $\underline{L}$ (as in Definition 1.1) on a closed surface $\Sigma_{g}$ does not vanish. Indeed, one can write $\underline{L}=\underline{L}^{\prime}\cup\alpha_{1}\cup...\alpha_{g}$ and decompose $\Sigma_{g}$ as a connected sum of a sphere containing $\underline{L}^{\prime}$ and $g$ tori each containing a circle $\alpha_{i}$. $\underline{L}^{\prime}$ is then a monotone link on the sphere, and by [CGHMSS22] its homology $HF(\underline{L}^{\prime},\underline{L}^{\prime})$ does not vanish. Applying the Künneth isomorphism $g$ times, we get that $HF(\underline{L},\underline{L})$ does not vanish either. Finally, by invariance of the homology, it also does not vanish for any choice of Hamiltonian $H$. Alternatively, one could try to directly prove the non-vanishing of the homology as in [CGHMSS22] by computing the potential of $\text{Sym}(\underline{L})$ and showing it admits a non-degenerate critical point. This potential is a map $W:H^{1}(\text{Sym}(\underline{L}),\mathbb{C}^{*})\to\mathbb{C}$ whose definition involves counting holomorphic discs of Maslov index 2 with boundary on $\text{Sym}\underline{L}$ (see [FOOO] for details about the theory of the potential).
3 The braid type function
3.1 The definition
In the introduction it was claimed that $b$ was a well defined function from $\text{Ham}_{c}(\Sigma_{g,p})$. To prove it, we need to remark that $b$ does not depend on the Hamiltonian isotopy one chooses. This is true if, for instance, $\pi_{1}(\text{Ham}_{c}(\Sigma_{g,p}))=0$. In fact, since in such a case any two Hamiltonian isotopies are homotopic relative endpoints between isotopies, we may deform one Hamiltonian path into the other keeping strands from crossing. Such a deformation then provides a braid isotopy between the images of the two Hamiltonian isotopies, and $b$ is well defined. We then just have to show that $\pi_{1}(\text{Ham}_{c}(\Sigma_{g,p}))=0$: this fact is standard, but as we could not find a reference, we decided to include the following proof.
Lemma 3.1.
If $\Sigma_{g,p}\neq\mathbb{S}^{2}$, $\text{Ham}_{c}(\Sigma_{g,p})$ is simply connected.
Proof.
Let $B=\partial\Sigma_{g,p}$: it is a disjoint union of $p$ circles. The embedding $B\hookrightarrow\Sigma_{g,p}$ provides a fibration
$\mathrm{Diff}_{0}(\Sigma_{g,p},B)\hookrightarrow\mathrm{Diff}_{0}(\Sigma_{g,p})\rightarrow\mathrm{Diff}_{0}(B)$ (8)
where the fibre is the group of diffeomorphisms of $\Sigma_{g,p}$ inducing the identity on the boundary and isotopic to the identity, the total space is the connected component of the identity of the diffeomorphisms of $\Sigma_{g,p}$, and the base is the connected component of the identity in the group of diffeomorphisms of a disjoint union of circles.
Let us consider the long exact sequence of the fibration in (8):
$\pi_{2}(\mathrm{Diff}_{0}(B))\rightarrow\pi_{1}(\mathrm{Diff}_{0}(\Sigma_{g,p},B))\rightarrow\pi_{1}(\mathrm{Diff}_{0}(\Sigma_{g,p}))\rightarrow\pi_{1}(\mathrm{Diff}_{0}(B))$ (9)
The term $\pi_{2}(\mathrm{Diff}_{0}(B))$ is always 0, since it is a power of $\pi_{2}(\mathrm{Diff}_{0}(\mathbb{S}^{1}))$ and this is trivial ($\mathrm{Diff}_{0}(\mathbb{S}^{1})$ has the homotopy type of the circle itself). This proves that the second arrow in the exact sequence (9) is always an injection. Now, if $g>1$ or $g=0,p\geq 3$, [Gramain, Theorem 1] shows that $\mathrm{Diff}_{0}(\Sigma_{g,p})$ is contractible: in such a case we deduce that $\pi_{1}(\mathrm{Diff}_{0}(\Sigma_{g,p},B))=0$. We want to prove that in this case $\mathrm{Diff}_{0}(\Sigma_{g,p},B)$ and $\text{Ham}_{c}(\Sigma_{g,p})$ have isomorphic fundamental groups.
Let $\text{Symp}_{0}(\Sigma_{g,p},B)$ be the group of symplectic diffeomorphisms of $\Sigma_{g,p}$ inducing the identity on the boundary and isotopic to the identity through symplectic diffeomorphisms with the same property. By [ban74] (generalising a result of Moser [mos65]) $\text{Symp}_{0}(\Sigma_{g,p},B)$ is a deformation retract of $\mathrm{Diff}_{0}(\Sigma_{g,p},B)$, and in particular
$\pi_{1}(\text{Symp}_{0}(\Sigma_{g,p},B))\cong\pi_{1}(\mathrm{Diff}_{0}(\Sigma_{g,p},B))$
Fix a decreasing sequence of open collar neighbourhoods of the boundary, call them $(U_{n})_{n\geq 1}$: they satisfy the property
$\bigcap_{n\geq 1}U_{n}=B$ (10)
Define by $\text{Symp}_{U_{n}}(\Sigma_{g,p})$ the group of symplectic diffeomorphisms of $\Sigma_{g,p}$ which are supported in $\Sigma_{g,p}\setminus U_{n}$ and are isotopic to the identity via such diffeomorphisms.
Clearly,
$\text{Symp}_{U_{n}}(\Sigma_{g,p})\subset\text{Symp}_{U_{n+1}}(\Sigma_{g,p}),\,\,\,\bigcup_{n\geq 1}\text{Symp}_{U_{n}}(\Sigma_{g,p})=\text{Symp}_{c}(\Sigma_{g,p})$ (11)
and both these conditions together imply that
$\pi_{1}(\text{Symp}_{c}(\Sigma_{g,p}))=\lim_{n}\pi_{1}(\text{Symp}_{U_{n}}(\Sigma_{g,p}))$ (12)
We shall now prove that $\pi_{1}(\text{Symp}_{c}(\Sigma_{g,p}))$ injects into $\pi_{1}(\text{Symp}_{0}(\Sigma_{g,p},B))=0$ via the map induced by the inclusion. From this it is possible to deduce that $\pi_{1}(\text{Ham}_{c}(\Sigma_{g,p}))=0$, since the map:
$\pi_{1}(\text{Ham}_{c}(\Sigma_{g,p}))\rightarrow\pi_{1}(\text{Symp}_{c}(\Sigma_{g,p}))$
defined by the inclusion is in fact an injection (see [mcDSal16, Proposition 10.2.13]).
Let therefore $t\mapsto\varphi_{t}\in\text{Symp}_{c}(\Sigma_{g,p})$ be a loop based at the identity which becomes contractible once seen as a representative in $\pi_{1}(\text{Symp}_{0}(\Sigma_{g,p},B))$. There is therefore a homotopy of symplectic isotopies fixing the boundary pointwise shrinking $t\mapsto\varphi_{t}$ to the constant path at the identity:
$\psi:[0,1]_{s}\times[0,1]_{t}\rightarrow\text{Symp}_{0}(\Sigma_{g,p},B),$
$\,\,\,\psi(0,t)=\varphi_{t},\psi(1,t)=\mathrm{Id},\psi(s,0)=\psi(s,1)=\mathrm{Id}$
The goal now is to modify this isotopy to a new one, in area-preserving diffeomorphisms which fix a small enough collar neighbourhood of the boundary. Since the support of $\varphi_{t}$ is compact in $\Sigma_{g,p}\setminus B$ for all $t$, there exists an $n$ big enough such that the support of $\varphi_{t}$ is contained in $\Sigma_{g,p}\setminus U_{n}$ for all $t$.
Since we know that $\text{Symp}_{0}(\Sigma_{g,p},B)$ is a deformation retract of $\text{Diff}_{0}(\Sigma_{g,p},B)$, we may homotope $\psi$ to another homotopy $\psi^{\prime}$, this time via diffeomorphisms fixing $U_{n}$ pointwise (but not necessarily area-preserving), from the identity to itself. We now consider $\psi^{\prime}$ as a homotopy in diffeomorphisms of $\overline{\Sigma_{g,p}\setminus U_{n}}$: applying Moser’s result again we find a homotopy in $\text{Symp}_{U_{n}}(\Sigma_{g,p})$, call it $\psi^{\prime\prime}$, between the symplectic loop $t\mapsto\varphi_{t}$ and the constant loop. Since $\text{Symp}_{U_{n}}(\Sigma_{g,p})\subset\text{Symp}_{c}(\Sigma_{g,p})$, the existence of $\psi^{\prime\prime}$ proves that we have a sequence of group morphisms
$\pi_{1}(\text{Ham}_{c}(\Sigma_{g,p}))\hookrightarrow\pi_{1}(\text{Symp}_{c}(\Sigma_{g,p}))\hookrightarrow\pi_{1}(\text{Symp}_{0}(\Sigma_{g,p},B))\cong\pi_{1}(\mathrm{Diff}_{0}(\Sigma_{g,p},B))$
and the rightmost group is 0 whenever $g>1$ or $g=0,p\geq 3$ by Gramain’s result. We have thus proved that $\text{Ham}_{c}(\Sigma_{g,p})$ is simply connected under the above topological assumptions.
We are now left with the cases of the disc $\mathbb{D}$ and cylinder $Z$ to consider. In these two cases [Gramain] shows that
$\text{Diff}_{0}(\mathbb{D},\partial\mathbb{D})\sim\text{Diff}_{0}(Z,\partial Z)\sim SO(2,\mathbb{R})$
where $\sim$ denotes homotopy equivalence, the diffeomorphisms induce the identity on the boundary, and $O(2,\mathbb{R})$ is the real orthogonal group of rank 2. Let us again adopt the notation from above: $\Sigma_{g,p}$ is our surface (disc or cylinder), and $B$ its boundary. We have again an exact sequence
$0\rightarrow\pi_{1}(\mathrm{Diff}_{0}(\Sigma_{g,p},B))\rightarrow\pi_{1}(\mathrm{Diff}_{0}(\Sigma_{g,p}))\rightarrow\pi_{1}(\mathrm{Diff}_{0}(B))$
but here the third group is not trivial. We need thus to show that the image of the second arrow is 0.
If $\Sigma_{g,p}=\mathbb{D}$, then $SO_{2}(\mathbb{R})$ is a subgroup of $\mathrm{Diff}_{0}(\Sigma_{g,p})$, and is in fact its deformation retract. A generator of $\pi_{1}(SO_{2}(\mathbb{R}))$ is given by a full rotation, which is mapped to the generator of $\pi_{1}(\text{Diff}_{0}(\partial\mathbb{D}))$: the third arrow is an injection in the case of the disc, so that the map
$\pi_{1}(\mathrm{Diff}_{0}(\Sigma_{g,p},B))\rightarrow\pi_{1}(\mathrm{Diff}_{0}(\Sigma_{g,p}))$
from above is 0 as claimed.
If we examine $\Sigma_{0,2}=Z$ instead, the argument is similar: the generator of $\pi_{1}(\text{Diff}_{0}(Z))$ is the full rotation, which is mapped to a non-zero element in $\pi_{1}(\text{Diff}_{0}(\partial Z))$ (not a generator in this case, of course). The third arrow is an injection in this case as well.
Summing up, in the cases $g=0,p\in\{1,2\}$ we still have $\pi_{1}(\text{Diff}_{0}(\Sigma_{g,p},B))=0$. The rest of the arguments above did not depend on the actual surface we worked on, and they carry over to this context: we infer that
$\pi_{1}(\text{Ham}_{c}(\mathbb{D}))\cong\pi_{1}(\text{Ham}_{c}(Z))=0$
Remark 3.2.
In fact, one should be able to deduce from the proof that $\text{Ham}_{c}(\Sigma_{g,p})$ is contractible whenever $\text{Diff}_{0}(\Sigma_{g,p},B)$ is. One may apply [pal66, Theorem 15] and use it in combination with the fact that the homotopy groups of an increasing union of topological spaces is the limit of homotopy groups of said subspaces (easy consequence of the compactness of spheres).
3.2 The image
Let $\Sigma_{g,p}$ be a Riemann surface with genus $g$ and $p$ punctures. Bellingeri in [bel04] describes via generators and relations the braid group with $n$ strands on $\Sigma_{g}$, for which we $\mathcal{B}_{n,g,p}$.
Such group has $n+2g+p-2$ generators, denoted by
$\sigma_{1},\dots,\sigma_{n-1},a_{1},\dots,a_{g},b_{1},\dots,b_{g},z_{1},\dots,z_{p-1}$ (13)
The $\sigma_{i}$ here represent a clockwise half-twist between $P_{i}$ and $P_{i+1}$ (while the other strands are fixed), the $a_{i}$ represent a non-contractible loop on a generator of the homology for the closed surface of genus $g$ based at the first base point, the $b_{i}$ is an analogous collection for the other $g$ generators, and $z_{i}$ stands for a non-contractible loop based at the first base point and going around the $i-th$ puncture. These generators are depicted in Figure 1.
$P_{1}$
$P_{n}$
$P_{1}$
$P_{j+1}$
$P_{n}$
$P_{j}$
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$P_{1}$
$P_{j}$
$P_{j+1}$
$P_{n}$
$P_{1}$
$P_{j}$
$P_{j+1}$
$P_{n}$
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$\zeta_{l}$
Figure 1: The generators of the braid group $\mathcal{B}_{n,g,p}$. In clockwise order, starting from top-left: $\sigma_{j}$, $a_{i}$, $z_{l}$ and $b_{i}$ We draw a fundamental domain of the surface, and all base points which are not endpoints of the green paths are to be thought of as constant paths.
There are then 8 groups of relations, we refer to [bel04, Theorem 1.1] for the precise definitions.
We aim to describe the braid type of a Hamiltonian diffeomorphism preserving a given premonotone Lagrangian configuration $\underline{L}$. Consider the subgroup of $\mathcal{B}_{k,g,p}$ generated by ${(\sigma_{i})}_{i=1,\dots,k-1}$, ${(a_{j})}_{j=1,\dots,g}$, ${(b_{j}^{-1}a_{j}b_{j})}_{j=1,\dots,g}$ and ${(z_{l})}_{l=1,\dots,p-1}$ only (no $b_{j}$ alone is present), with the restrictions of Bellingeri’s relations. For clarity, the $\sigma_{i}$ we consider are those describing exchanges between contractible components of the link only. Denote this subgroup by $\mathcal{B}_{\underline{L}}.$
Remark 3.3.
Even though we have $p$ punctures, the associated generators are only $p-1$. As remarked in [bel04], one may express a single loop $z_{p}$ around the last puncture using the relation
$[a_{1},b_{1}^{-1}]\cdots[a_{g},b_{g}^{-1}]=\sigma_{1}\sigma_{2}\cdots\sigma_{k-1}^{2}\cdots\sigma_{2}\sigma_{1}z_{1}^{-1}\cdots z_{p}^{-1}$
Lemma 3.4.
Let us consider the group homomorphism
$b:\text{Ham}_{\underline{L},c}(\Sigma_{g})\rightarrow\mathcal{B}_{k,g,p}$ (14)
Its image is precisely $\mathcal{B}_{\underline{L}}$.
Proof.
We are going to describe the group $\mathcal{B}_{\underline{L}}$ as the fundamental group of the $k$-th configuration space of the surface $\tilde{\Sigma}_{g,p}$, which we obtain by removing tubular neighbourhoods of the $g$ $\beta$ curves from $\Sigma_{g,p}$. Now, $\tilde{\Sigma}_{g,p}$ is a punctured sphere, with $2g+p$ punctures. Its braid group is described in [beFu04, Theorem 2.1]: it has the $k-1$ generators corresponding to moves which take place in a disc on the surface, and $p+2g-1$ corresponding to non-contractible loops around punctures.
Let us show that we may identify the image of $b$ with this fundamental group. Let $\varphi\in\text{Ham}_{\underline{L},c}(\Sigma_{g,p})$, and let $L_{j}$ be any non contractible component of $\underline{L}$. Since it cannot be displaced by any Hamiltonian diffeomorphism, if $(\phi^{t})$ is any homotopy between the identity of Ham and $\varphi$, $L_{j}$ satisfies
$L_{j}\cap\phi^{t}(L_{j})\neq\emptyset$
This implies that there is a loop in $\text{Ham}_{c}(\Sigma_{g,p})$ based at the identity, $t\mapsto\psi^{t}$, such that for all $t$ and for any non contractible component $L_{j}$ one has
$\psi^{t}\phi^{t}(L_{j})=L_{j}$
Since the circles cannot intersect during this isotopy, and since $b$ does not depend on the particular isotopy one chooses without loss of generality (up to shrinking the tubular neighbourhoods we take off to produce $\tilde{\Sigma}_{g,p}$) the braid $b(\varphi)$ is represented by a braid entirely contained in $\tilde{\Sigma}_{g,p}$. Furthermore, choosing such an isotopy makes it clear that no contractible circle may wind about the base point on a non-contractible component: this bounds the number of $\sigma$-generators to $k-1$ (only exchanges between contractible circles are allowed). All of this proves that the image of $b$ is contained in $\pi_{1}(\mathrm{Conf}^{k}(\tilde{\Sigma}_{g,p}))$, which is what we wanted. To prove equality, we just need to be able to represent all generators of $\mathcal{B}_{\underline{L},k}$ as images of $b$. This is readily done: given any $\sigma_{i}$, we may consider a disc containing $L_{1}$ and $L_{i}$ and apply a half rotation. In the case of any $z_{i}$ (loop around a puncture) we may use an annulus with centre at the puncture, and a full rotation on it.
We end this proof showing that the generators above, elements in $\mathcal{B}_{k,g,p}$, are in fact the images of the generators of $\pi_{1}(\mathrm{Conf}^{k}(\tilde{\Sigma}_{g,p}))$ under the inclusion morphism. In fact, it is easy to see that the $\sigma_{i}$’s in $\mathcal{B}_{k,g,p}$ correspond to the half-twists in $\pi_{1}(\mathrm{Conf}^{k}(\tilde{\Sigma}_{g,p}))$, and that the turns around the punctures in $\tilde{\Sigma}_{g,p}$ are sent by the inclusion to turns around the punctures in $\Sigma_{g,p}$ (the $z_{j}$’s) and turns around the handles (the $a_{i}$’s and $b_{i}^{-1}a_{i}b_{i}$’s) (see Figure 2 for a picture of $b_{i}^{-1}a_{i}b_{i}$).
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$L_{k+i}$
$L_{1}$
$L_{k}$
$b_{i}^{-1}a_{i}b_{i}$
Figure 2: The generator $b_{i}^{-1}a_{i}b_{i}$
Remark 3.5.
If $g=0,p=1$ this Lemma proves surjectivity of the braid type function on the disc.
4 Proof of the main result
4.1 Intersections in the symmetric product
We define elementary intersections between two strands: they are local descriptions for intersections counting towards $u\cdot\Delta$ which do not lift along the projection to the symmetric product. The goal of this section is to prove that elementary intersections are transverse in the symmetric product, and may thus be used to compute the intersection product with the diagonal.
First off, let us reduce the computation of the intersection product to counting the intersections between two strands.
Let $u:[0,1]\times[0,1]\rightarrow\text{Sym}^{k+g}(\Sigma_{g,p})$ be a capping for an intersection point between the $\text{Sym}^{k+g}\underline{L}$ and $\text{Sym}^{k+g}\varphi(\text{Sym}^{k+g}\underline{L})$. Assume that for some $(s_{0},t_{0})$, $u(s_{0},t_{0})\in\Delta$. Generically, $u(s_{0},t_{0})$ belongs in the top stratum of the diagonal, which means that considering the path in the symmetric product
$t\mapsto u(s_{0},t)$
as a collection of $k+g$ curves, at $t_{0}$ exactly two of them coincide (and there are no other intersections at $t_{0}$ or different $t$-we may assume this generically). Denote these two paths by $u_{2}:[0,1]\times[0,1]\rightarrow\text{Sym}^{2}\Sigma_{g}$, and the 2-dimensional diagonal by $\Delta_{2}$.
Lemma 4.1 (Locality of the intersection problem).
In the setting as above, the sign of the intersection at $(s_{0},t_{0})$ between $u$ and $\Delta$ is the same of the intersection at $(s_{0},t_{0})$ between $u_{2}$ and $\Delta_{2}$.
Proof.
Let us now consider a local chart of $\text{Sym}^{k+g}(\Sigma_{g})$ adapted to $\Delta$ at the intersection $u(s_{0},t_{0})=[x_{1},x_{1},x_{3},\dots,x_{k+g}]\in\Delta$. Consider the function $Sym^{2}(\Sigma_{g})\rightarrow\text{Sym}^{k+g}(\Sigma_{g})$ defined via
$[y_{1},y_{2}]\mapsto[y_{1},y_{2},x_{3},\dots,x_{k+g}]$ (15)
This map is a complex embedding: it is clearly a continuous injective maps between compact topological spaces, hence a homeomorphism on its image, and its differential is injective. To prove this last point, we consider local charts given by symmetric polynomials. In formulas,
$\text{Sym}^{2}(\mathbb{C})\cong\mathbb{C}^{2},[y_{1},y_{2}]\mapsto(y_{1}+y_{2},y_{1}y_{2})$ (16)
$\text{Sym}^{k+g}(\mathbb{C})\cong\mathbb{C}^{k+g},[x_{1},\dots,x_{k+g}]\mapsto(e_{1}(x_{1},\dots,x_{k+g}),\dots,e_{k+g}(x_{1},\dots,x_{k+g}))$ (17)
In the equation above, $e_{i}$ is the $i$-th symmetric polynomials on $k+g$ variables,
$e_{i}(x_{1},\dots,x_{k+g}):=\sum_{1\leq l_{1}<\dots<l_{i}\leq k+g}\prod_{j=1}^{i}x_{l_{j}}$ (18)
In particular, $e_{1}$ and $e_{2}$ on two variables are the expressions appearing in (16). To prove injectivity of the differential, we just have to compute four derivatives. When restricting to the image of the topological embedding,
$e_{1}(y_{1},y_{2},x_{3},\dots,x_{k})=y_{1}+y_{2}+x_{3}+\cdots+x_{k+g}$ (19)
$e_{2}(y_{1},y_{2},x_{3},\dots,x_{k+g})=(y_{1}+y_{2})(x_{3}+\cdots+x_{k+g})+y_{1}y_{2}+x_{3}x_{4}+\cdots x_{k+g-1}x_{k+g}$ (20)
One can see that the topological embedding is in fact complex. Now, in local charts,
$\partial_{y_{1}+y_{2}}e_{1}=1,\,\,\partial_{y_{1}y_{2}}e_{1}=0,\,\,\partial_{y_{1}+y_{2}}e_{2}=x_{3}+\cdots+x_{k+g},\,\,\partial_{y_{1}y_{2}}e_{2}=1$ (21)
that which proves injectivity of the differential. Now, we see $\text{Sym}^{2}(\Sigma_{g})$ as a complex submanifold of $\text{Sym}^{k+g}(\Sigma_{g})$, and remark that the embedding maps $\Delta_{2}$ into $\Delta$.
Assume now that the $k+g-2$ strands which do not intersect do not move on an interval $(s_{0}-\varepsilon,s_{0}+\varepsilon)$ (we do not lose generality since we are just reparametrising the capping). This condition implies that $Im(d_{(s_{0},t_{0})}u)\subset T_{(u(s_{0},t_{0}))}\text{Sym}^{2}(\Sigma_{g,p})$ (it is implicit in the notation that $u(s_{0},t_{0})\in\Delta_{2}$). Now, choose a complex chart around $u(s_{0},t_{0})$ adapted to $\text{Sym}^{2}(\Sigma_{g,p})$: such a chart maps $\text{Sym}^{2}(\Sigma_{g,p})$ locally to $\mathbb{C}^{2}\times\{0\}^{k+g-2}$ and $T_{u(s_{0},t_{0})}\Delta$ also splits:
$T_{u(s_{0},t_{0})}\Delta\cong V\oplus\mathbb{C}^{k+g-1}$
where $V$ is a complex line in $\mathbb{C}^{2}$ (we have $\mathbb{C}^{k+g-1}$ as second factor because the other $k+g-1$ coordinates may change as they wish, it won’t change the fact that the point is in the diagonal because of the first two coordinates). Therefore, in order to compute the sign of the intersection, we are led to consider the determinant of the real matrix
$M:=\begin{pmatrix}v&w&x&y&0&\cdots&\cdots&\cdots&0\\ 0&\cdots&\cdots&0&1&0&\cdots&\cdots&0\\ \vdots&&&\vdots&0&1&0&\cdots&0\\ \vdots&&&\vdots&\vdots&&\ddots&\ddots&\vdots\\ \vdots&&&\vdots&\vdots&&&\ddots&0\\ 0&\cdots&\cdots&0&0&\cdots&\cdots&0&1\\ \end{pmatrix}$ (22)
which is a block matrix. Here $v,w\in\mathbb{R}^{4}$ are a real basis of $Im(d_{(s_{0},t_{0})}u)$, and $x,y\in\mathbb{R}^{4}$ are similarly a real basis of $V$. Clearly
$\det(M)=\det(v,w,x,y)$ (23)
which is what we wanted to show. ∎
Definition 4.1.
Let $(s,t)\in[0,1]^{2}$. We say that a capping $u$ has an elementary intersection at $(s,t)$ with $\Delta$ if $u(s,t)\in\Delta$ and if there exist charts as those defined in Lemma 4.1 around $u(s,t)$ such that $u_{2}$ coincides with the roots of the complex polynomial
$X^{2}-(s+it)$
Lemma 4.2.
Elementary intersections between two strands are transverse.
Proof.
We will start by considering a homotopy $\gamma:[-1,1]_{s}\times[-1,1]_{t}\rightarrow\text{Sym}^{2}(\mathbb{C})$ between two braids $\gamma_{-}$ and $\gamma_{+}$. There is a diffeomorphism $\phi:\mathbb{C}^{2}\rightarrow\text{Sym}^{2}(\mathbb{C})$ which maps $(a,b)\in\mathbb{C}^{2}$ to the pair of roots of the degree 2 polynomial $X^{2}-aX+b$ (its inverse being given by $\phi^{-1}([x_{-},x_{+}])=(x_{-}+x_{+},x_{-}x_{+})$). Through this diffeomorphism, the diagonal $\Delta\subset\text{Sym}^{2}(\mathbb{C})$ corresponds to the set
$\phi^{-1}(\Delta)=\left\{(2x,x^{2}),x\in\mathbb{C}\right\}=\left\{\left(a,\frac{a^{2}}{4}\right),a\in\mathbb{C}\right\}$
whose tangent space at $(a,\frac{a^{2}}{4})$ is the complex vector subspace generated by $(1,\frac{a}{2})$, i.e. the real vector subspace generated by $(1,\frac{a}{2})$ and $(i,\frac{ia}{2})$. Let $(a(s,t),b(s,t)):=\phi^{-1}(\gamma(s,t))$. Assume that $\gamma$ intersects the diagonal at $(s,t)=(0,0)$. Then, the intersection is transverse if and only if the vectors $(1,\frac{a(0,0)}{2})$, $(i,\frac{ia(0,0)}{2})$, $(\partial_{s}a(0,0),\partial_{s}b(0,0))$ and $(\partial_{t}a(0,0),\partial_{t}b(0,0))$ generate the whole space $\mathbb{C}^{2}$ as a real vector space.
In particular, consider the case where $\gamma(s,t)$ is the pair of square roots of $s+it$. Then, $\gamma$ intersects the diagonal at $(0,0)$, and $(a(s,t),b(s,t))=(0,-s-it)$. Therefore, we get that the intersection is transverse if and only if the vectors $(1,0),(i,0),(0,-1),(0,-i)$ generate $\mathbb{C}^{2}$. Since this is true, we have transversality of the intersection.
$t$
$t\mapsto\gamma(-0.01,t)$
$t\mapsto\gamma(0,t)$
$t\mapsto\gamma(0.01,t)$
Figure 3: The homotopy $\gamma(s,t)$
$s$
$t\mapsto\gamma(-0.01,t)$
$t\mapsto\gamma(0,t)$
$t\mapsto\gamma(-0.01,t)$
Figure 4: The homotopy $\gamma(s,t)$ as seen on $\mathbb{C}$.
Remark 4.3.
From the proof we see that the sign of the intersection we studied above as local model is positive.
Remark 4.4.
The homotopy $s+it\mapsto\sqrt{(s+it)}\in\text{Sym}^{2}\mathbb{C}$ does not lift to a function to $\mathbb{C}^{2}$. If we, in more generality, suppose that an intersection is transverse, the capping cannot be lifted to $\mathbb{C}^{2}$. Assume in fact that an intersection counting towards $[u]\cdot\Delta$ does lift to $\mathbb{C}^{2}$ (we take $k=2$ in light of Lemma 4.1). Then it cannot be transverse: up to time-translation, assume the intersection appears at $(s,t)=(0,0)$. Define
$(\gamma_{1},\gamma_{2}):(-\varepsilon,\varepsilon)\times(-\varepsilon,\varepsilon)\rightarrow\mathbb{C}^{2}$
such that if $\pi:\mathbb{C}^{2}\rightarrow\text{Sym}^{2}\mathbb{C}$ is the quotient projection, we locally have
$\pi\circ(\gamma_{1},\gamma_{2})=u$
Then $(\gamma_{1},\gamma_{2})(0,0)\in\pi^{-1}(\Delta)$. The differential of $\pi$ being 0 there, so is the differential of $u$. This proves that the intersection of $u$ with $\Delta$ at $(0,0)$ is not transverse.
4.2 Construction of the Hofer-Lipschitz function
Modulo modifications due to the new context with just $g$ non-contractible components, the proof in [M] goes through, we only need some minor modifications.
Let $\zeta_{1},...,\zeta_{p}$ be the boundary components of $\Sigma_{g,p}$. Gluing discs of area $s_{i,j}$ along $\zeta_{j}$ for $i=1,2$, $j=1,...,p$ gives rise to two embeddings of $\Sigma_{g,p}$ into a closed surface of area $1+s_{i}$, $i=1,2$, denoted $\Sigma_{g}(1+s_{i})$, where
$s_{i}=s_{i,1}+...+s_{i,p}\in(0,(k+1)A-1)$
In the following, by $v_{i}$ we denote the vector
$v_{i}:=(s_{i,1},\dots,s_{i,p})$
and the set of possible values $v_{1},v_{2}$ may take is
$V:=\{(s_{1},...,s_{p})\in(\mathbb{R}_{\geq 0})^{p}|s_{1}+...+s_{p}\leq(k+1)A-1\}$
Choose a diffeomorphism between the two closed surfaces: up to changing the complex structure on one of them it becomes a biholomorphism. We also assume this diffeomorphism to be the identity between small neighbourhoods of $\Sigma_{g,p}$ in $\Sigma_{g}(1+s_{i})$.
This operation induces a bijective correspondence $\Phi$ between capping classes: it is the content of [M, Lemma 3.2] which still holds here (we need for this step the generalisation of monotonicity as proved in Section 2.2). In our setting we may also prove, (see [M, Lemmata 3.3, 3.4] ), that gluing discs of these areas gives an isomorphism at the chain level of Floer complexes, and that this isomorphism commutes with the PSS maps. In [M] the conventions are slightly different, and for the differential one counts holomorphic strips, while here the strips have to solve the Floer equation. This does not represent a problem in our context: since the Hamiltonian is 0 on the discs we glue, the biholomorphism between surfaces $\Sigma_{g}(1+s_{i})$ still defines a chain map since it sends the Hamiltonian vector field on $\Sigma_{g}(1+s_{1})$ to the one on $\Sigma_{g}(1+s_{2})$.
We still need to compute the action shift induced by the change in the areas: we do this in the following Lemma. Before stating it, we recall the basic notations and notions the computation involves.
Let $\varphi\in\text{Ham}_{\underline{L},c}(\Sigma_{g,p})$, and let $H$ be a Hamiltonian generating $\varphi$, supported in the interior of $\Sigma_{g,p}$. Let $\underline{L}_{v_{i}}$ be the image of $\underline{L}$ inside $\Sigma_{g}(1+s_{i})$, $i=1,2$, and $H_{v_{i}}$ the extension by 0 of $H$ in the closed surfaces. After a small perturbation of $H$, we can assume that the Heegaard-Floer complexes for $(\underline{L}_{v_{i}},H_{v_{i}})$, $i=1,2$ are well defined. Using the bijective correspondence between cappings above, $[\hat{y}_{1}]$ corresponds to a generator $[\hat{y}_{2}]:=\Phi([\hat{y}_{1}])$ of $CF^{*}(\text{Sym}^{k+g}H_{v_{2}},\text{Sym}\underline{L}_{v_{2}})$.
Lemma 4.5.
The difference of action $\mathcal{A}_{H_{v_{1}}}([\hat{y}_{1}])-\mathcal{A}_{H_{v_{2}}}([\hat{y}_{2}])$ does not depend on the choice of the generator $[\hat{y}_{1}]$.
Proof.
Start by fixing a generator $[\hat{y}_{1}]$. Since $H_{v_{i}}$, $i=1,2$ are supported in $\Sigma_{g,p}$, we get that $y_{i}$ are paths in $\text{Sym}^{k+g}(\Sigma_{g,p})$. The biholomorphism between $\Sigma_{g}(1+s_{1})$ and $\Sigma_{g}(1+s_{2})$ restricts to the identity on $\Sigma_{g,p}$, so the two paths $y_{1}$ and $y_{2}$ coincide, and $H_{v_{1}}|_{y_{1}}=H_{v_{2}}|_{y_{2}}=H|_{y_{1}}$. Therefore, using the formula for the action, the terms containing the integral of the Hamiltonian will cancel out, leaving us with:
$\mathcal{A}_{H_{v_{1}}}([\hat{y}_{1}])-\mathcal{A}_{H_{v_{2}}}([\hat{y}_{2}])=\eta_{s_{2}}[\hat{y}_{2}]\cdot\Delta-\eta_{s_{1}}[\hat{y}_{1}]\cdot\Delta+\omega_{s_{2}}([\hat{y}_{2}])-\omega_{s_{1}}([\hat{y}_{1}])$
We first show that this quantity does not depend on the choice of capping for the path $y_{1}$. Let $[\hat{y}^{\prime}_{1}]$ be another capping. We want to show that
$\eta_{s_{2}}([\hat{y}_{2}]\#\Phi[\hat{y}^{\prime}_{1}]^{-1})\cdot\Delta+$ $\omega_{s_{2}}([\hat{y}_{2}]\#\Phi[\hat{y}^{\prime}_{1}]^{-1})-$
-- $\left(\eta_{s_{1}}([\hat{y}_{1}]\#[\hat{y}^{\prime}_{1}]^{-1})\cdot\Delta+\omega_{s_{1}}([\hat{y}_{1}]\#[\hat{y}^{\prime}_{1}]^{-1})\right)=0$
Since $[\hat{y}_{1}]\#[\hat{y}^{\prime}_{1}]^{-1}$ and $\Phi[\hat{y}_{1}]\#\Phi[\hat{y}^{\prime}_{1}]^{-1}$ are homotopies between $x$ and itself, it is in fact a disc with boundary on $\text{Sym}\underline{L}$, so by monotonicity (Proposition 2.1),
$\eta_{s_{1}}([\hat{y}_{1}]\#[\hat{y}^{\prime}_{1}]^{-1})\cdot\Delta+\omega_{s_{1}}([\hat{y}_{1}]\#[\hat{y}^{\prime}_{1}]^{-1})=\frac{\lambda}{2}\mu([\hat{y}_{1}]\#[\hat{y}^{\prime}_{1}]^{-1})$
$\eta_{s_{2}}($ $\Phi[\hat{y}_{1}]\#\Phi[\hat{y}^{\prime}_{1}]^{-1})\cdot\Delta+\omega_{s_{2}}(\Phi[\hat{y}_{1}]\#\Phi[\hat{y}^{\prime}_{1}]^{-1})=\frac{\lambda}{2}\mu(\Phi[\hat{y}_{1}]\#\Phi[\hat{y}^{\prime}_{1}]^{-1})$
Since the Maslov index is preserved by the bijective correspondence, the difference above vanishes, which proves that the difference of action does not depend on the choice of capping. Now we show that it does not depend on the path $y_{1}$ either. If $y^{\prime}_{1}$ is another trajectory of $H_{v_{1}}$ between $\text{Sym}\underline{L}_{v_{1}}$ and itself, then we can define a capping for $y^{\prime}_{1}$ by choosing any capping for $y_{1}$ and concatenating it with any homotopy $[w_{1}]$ from $y_{1}$ to $y^{\prime}_{1}$. Since we showed that the action difference does not depend on the capping, it is enough to find a single homotopy $[w_{1}]$ for which
$\eta_{s_{2}}(\Phi[w_{1}])\cdot\Delta+\omega_{s_{2}}(\Phi[w_{1}])-\left(\eta_{s_{1}}([w_{1}])\cdot\Delta+\omega_{s_{1}}([w_{1}])\right)=0$
Such a homotopy exists: we can first find a homotopy sliding the end points of $y_{1}$ and $y_{1}^{\prime}$ inside the circles so that they coincide (and therefore define braids), and then we can take a homotopy inside $\text{Sym}^{k+g}(\Sigma_{g,p})$ that does not intersect $\Delta$, because $y_{1}$ and $y^{\prime}_{1}$ have the same braid type. ∎
The previous Lemma implies that for any $\varphi\in\text{Ham}_{\underline{L},c}(\Sigma_{g,p})$, and any choice of generator $[\hat{y}_{1}]$, we have $c_{\underline{L}_{v_{1}}}(\varphi)-c_{\underline{L}_{v_{2}}}(\varphi)=\frac{1}{k+g}(\mathcal{A}_{H_{v_{1}}}([\hat{y}_{1}])-\mathcal{A}_{H_{v_{2}}}([\hat{y}_{2}]))$ as in [M, Lemma 3.6].
Definition 4.2.
We define $f_{v_{1},v_{2}}:\text{Ham}_{c}(\Sigma_{g,p})\rightarrow\mathbb{R}$ as
$f_{v_{1},v_{2}}(\varphi):=c_{\underline{L}_{v_{1}}}(\varphi)-c_{\underline{L}_{v_{2}}}(\varphi)=\frac{1}{k+g}(\mathcal{A}_{H_{v_{1}}}([\hat{y}_{1}])-\mathcal{A}_{H_{v_{2}}}([\hat{y}_{2}]))$
Moreover, we have:
Lemma 4.6.
The map $f_{v_{1},v_{2}}:\text{Ham}_{\underline{L},c}(\Sigma_{g,p})\to\mathbb{R}$ is a group homomorphism, which factorises over $b:\text{Ham}_{\underline{L},c}(\Sigma_{g,p})\to\mathcal{B}_{\underline{L}}$, i.e. there exists a group homomorphism $\mathfrak{f}_{v_{1},v_{2}}:\mathcal{B}_{\underline{L}}\to\mathbb{R}$ such that $f_{v_{1},v_{2}}=\mathfrak{f}_{v_{1},v_{2}}\circ b$.
Proof.
Let $\varphi$, $\psi$ be in $\text{Ham}_{\underline{L},c}(\Sigma_{g,p})$. Let $H$ (resp. $H^{\prime}$) be a Hamiltonian supported inside $\Sigma_{g,p}$ generating $\varphi$ (resp. $\psi$). Let $[\hat{y}_{1}]$ (resp. $[\hat{y^{\prime}}_{1}]$) be a generator of $CF^{*}(\text{Sym}^{k+g}H_{v_{1}},\text{Sym}\underline{L}_{v_{1}})$ (resp. $CF^{*}(\text{Sym}^{k+g}H^{\prime}_{v_{1}},\text{Sym}\underline{L}_{v_{1}})$. Let $z:[0,1]\to\text{Sym}^{k+g}\Sigma_{g,p}$ be a trajectory of $H\#H^{\prime}$ from $\text{Sym}\underline{L}$ to itself. We want to construct a capping $\hat{z}_{1}:[0,1]\times[0,1]\to\text{Sym}^{k+g}\Sigma_{g}(1+s_{1})$ from $z$ to the constant path $x$, using the cappings $\hat{y}_{1}$ and $\hat{y^{\prime}}_{1}$. We define it as follows (see Fig. 5):
$\subset\text{Sym}(\underline{L})$
$\subset\text{Sym}(\underline{L})$
$\subset\text{Sym}(\underline{L})$
$z$
$x$
$z(1)$
$z(0)$
$\subset\text{Sym}^{k+g}(\Sigma_{g,p})\setminus\Delta$
$\hat{y}_{1}$
$\hat{y}^{\prime}_{1}$
$y^{\prime}$
$y$
Figure 5: The capping $\hat{z}_{1}$
  • On $[\frac{1}{2},1]\times[0,\frac{1}{5}]$:
    • $\hat{z}_{1}([\frac{1}{2},1]\times[0,\frac{1}{5}])\subset\text{Sym}\underline{L}_{v_{1}}$
    • for $0\leq t\leq\frac{1}{5}$, $\hat{z}_{1}(1,t)=x$
    • $\hat{z}_{1}(\frac{1}{2},0)=z(0)$
    • for $\frac{1}{2}\leq s\leq 1$, $\hat{z}_{1}(s,\frac{1}{5})=\hat{y}_{1}(2s-1,0)$
    (Such a homotopy exists because $\text{Sym}\underline{L}_{v_{1}}$ is path-connected, it is in fact isotopic to a Clifford torus, cf. [CGHMSS22])
  • On $[\frac{1}{2},1]\times[\frac{1}{5},\frac{2}{5}]$, $\hat{z}_{1}(s,t)=\hat{y}_{1}(2s-1,5t-1)$
  • On $[\frac{1}{2},1]\times[\frac{2}{5},\frac{3}{5}]$:
    • $\hat{z}_{1}([\frac{1}{2},1]\times[\frac{2}{5},\frac{3}{5}])\subset\text{Sym}\underline{L}_{v_{1}}$
    • for $\frac{2}{5}\leq t\leq\frac{3}{5}$, $\hat{z}_{1}(1,t)=x$
    • for $\frac{1}{2}\leq s\leq 1$, $\hat{z}_{1}(s,\frac{2}{5})=\hat{y}_{1}(2s-1,1)$
    • for $\frac{1}{2}\leq s\leq 1$, $\hat{z}_{1}(s,\frac{3}{5})=\hat{y^{\prime}}_{1}(2s-1,0)$
  • On $[\frac{1}{2},1]\times[\frac{3}{5},\frac{4}{5}]$, $\hat{z}_{1}(s,t)=\hat{y^{\prime}}_{1}(2s-1,5t-3)$
  • On $[\frac{1}{2},1]\times[\frac{4}{5},1]$:
    • $\hat{z}_{1}([\frac{1}{2},1]\times[\frac{4}{5},1])\subset\text{Sym}\underline{L}_{v_{1}}$
    • for $\frac{4}{5}\leq t\leq 1$, $\hat{z}_{1}(1,t)=x$
    • for $\frac{1}{2}\leq s\leq 1$, $\hat{z}_{1}(s,\frac{4}{5})=\hat{y^{\prime}}_{1}(2s-1,1)$
    • $\hat{z}_{1}(\frac{1}{2},1)=z(1)$
  • On $[0,\frac{1}{2}]\times[0,1]$, $\hat{z}_{1}$ is a homotopy of braids between $z$ and $\hat{z}_{1}(\frac{1}{2},\cdot)$, contained in $\text{Sym}^{k+g}(\Sigma_{g,p})$. Indeed, since $b$ is a group homomorphism, those two paths are isotopic as braids.
Now, the capping $\hat{z}_{1}$ consists of a concatenation of $\hat{y}_{1}$, $\hat{y^{\prime}}_{1}$, homotopies contained in $\text{Sym}\underline{L}_{v_{1}}$, and a braid isotopy contained in $\text{Sym}^{k+g}(\Sigma_{g,p})$. Homotopies contained in $\text{Sym}^{k+g}(\Sigma_{g,p})$ do not contribute to the difference of symplectic area, and as $\text{Sym}\underline{L}_{v_{1}}$ is away from the diagonal, and (by definition) a braid isotopy does not cross the diagonal, the only contribution to the difference of action comes from $\hat{y}_{1}$ and $\hat{y^{\prime}}_{1}$. Since the intersection number and the symplectic area are additive, we get that $f_{v_{1},v_{2}}(\varphi\psi)=f_{v_{1},v_{2}}(\varphi)+f_{v_{1},v_{2}}(\psi)$. Moreover, since braid isotopies do not contribute to the action difference, $f_{v_{1},v_{2}}(\varphi)$ only depends on the braid type of $\varphi$. ∎
Therefore, it is enough to compute the value of $\mathfrak{f}_{v_{1},v_{2}}$ on the generators of $\mathcal{B}_{\underline{L},k}$ to express $f_{v_{1},v_{2}}(\varphi)$ (granted that we know how to decompose the braid type of $\varphi$ as a product of generators).
Lemma 4.7.
The values of $\mathfrak{f}_{v_{1},v_{2}}$ are the following:
$\mathfrak{f}_{v_{1},v_{2}}(a_{i})=2\frac{\eta_{2}-\eta_{1}}{k+g}=-f_{v_{1},v_{2}}(b_{i}^{-1}a_{i}b_{i}),\,\,\,\mathfrak{f}_{v_{1},v_{2}}(\sigma_{j})=-\frac{\eta_{2}-\eta_{1}}{k+g},$
$\mathfrak{f}_{v_{1},v_{2}}(z_{j})=\frac{s_{2,j}-s_{1,j}}{k+g}=-2(\eta_{2}-\eta_{1})\frac{k+2g-1}{k+g}\frac{s_{2,j}-s_{1,j}}{s_{2}-s_{1}}$
Proof.
To do these computations, according to Lemma 4.5 it is enough to produce explicit homotopies between the reference path $x$ (which is a trivial braid) and the braid for which we know how to compute the action difference. For $a_{i}$, $b_{i}^{-1}a_{i}b_{i}$ and $\sigma_{j}$, we exhibit homotopies contained in $\Sigma_{g,p}$, so that the only contribution to the action difference comes from intersections with the diagonal. We also choose our homotopies so that the intersections with the diagonal are transverse. To be sure of their transversality, we make use the content of Section 4.1. We are going to show there that the intersections we count here are transverse, and give their signs.
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$L_{1}$
$L_{k}$
$L_{k+i}$
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$L_{1}$
$L_{k}$
$L_{k+i}$
$a_{i}$
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$L_{1}$
$L_{k}$
$L_{k+i}$
Figure 6: Homotopy between $a_{i}$ and the constant path
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$L_{1}$
$L_{k}$
$L_{k+i}$
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$L_{1}$
$L_{k}$
$L_{k+i}$
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$L_{1}$
$L_{k}$
$L_{k+i}$
Figure 7: Homotopy between $b_{i}^{-1}a_{i}b_{i}$ and the constant path
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$L_{k+i}$
$\sigma_{j}$
$L_{1}$
$L_{j}$
$L_{j+1}$
$L_{k}$
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$L_{k+i}$
$L_{1}$
$L_{j}$
$L_{j+1}$
$L_{k}$
$\alpha_{i}$
$\beta_{i}$
$\alpha_{i}$
$\beta_{i}$
$\zeta_{1}$
$\zeta_{p}$
$L_{k+i}$
$L_{1}$
$L_{j}$
$L_{j+1}$
$L_{k}$
Figure 8: Homotopy between $\sigma_{j}$ and the constant path
The result comes from counting such intersections (with sign) on Figures 6, 7 and 8. All the intersections appearing in those homotopies are modelled on the example of Figure 4, therefore we know they are transverse and can compute their signs.
For $z_{j}$, the braid becomes trivial after embedding $\Sigma_{g,p}$ into $\Sigma_{g}(1+s_{i})$, and we can choose an isotopy to the trivial braid (without crossing) which sweeps the disc glued to the boundary component $\zeta_{j}$. Therefore, the only contribution to the action difference comes from the difference of symplectic area between the two embeddings:
$\mathfrak{f}_{v_{1},v_{2}}(z_{j})=\frac{s_{2,j}-s_{1,j}}{k+g}$
By the monotonicity property applied to the only non-contractible connected component of $\Sigma_{g}(1+s_{i})$, we have
$A=1+s_{i}-kA+2\eta_{i}(k+2g-1)$
and therefore $\eta_{2}-\eta_{1}=\frac{s_{1}-s_{2}}{2(k+2g-1)}$, which implies that
$\mathfrak{f}_{v_{1},v_{2}}(z_{j})=-2(\eta_{2}-\eta_{1})\frac{k+2g-1}{k+g}\frac{s_{2,j}-s_{1,j}}{s_{2}-s_{1}}$
Remark 4.8.
With this last expression for $\mathfrak{f}_{v_{1},v_{2}}(z_{j})$, it is easy to check that the values of $\mathfrak{f}_{v_{1},v_{2}}$ are consistent with the relation of Remark 3.3.
Remark 4.9.
The reader might be surprised by the fact that
$\mathfrak{f}_{(v_{1},v_{2})}(b_{i}a_{i}b_{i}^{-1})\neq\mathfrak{f}_{(v_{1},v_{2})}(a_{i})$
Now, remark that the $\mathfrak{f}_{(v_{1},v_{2})}$ are not defined on the $b_{i}$, which indeed are not elements in $\mathcal{B}_{\underline{L}}$.
By the Hofer-Lipschitz property of link spectral invariants, we have that for all $\varphi\in\text{Ham}_{\underline{L},c}(\Sigma_{g,p})$, and for any choice of $(v_{1},v_{2})=(s_{i,j})_{i=1,2;1\leq j\leq p}\in V\times V$:
$\lVert\varphi\rVert\geq\frac{1}{2}\lvert\mathfrak{f}_{v_{1},v_{2}}(\varphi)\rvert$
To get the best estimate, we compute the maximum of $\lvert\mathfrak{f}_{(s_{i,j})}(\varphi)\rvert$ over the choice of $(v_{1},v_{2})$. We have
$(k+g)\mathfrak{f}_{v_{1},v_{2}}(\varphi)=\frac{s_{1}-s_{2}}{2(k+2g-1)}(2k_{gen}-k_{\sigma})+\sum\limits_{j=1}^{p-1}k_{j}(s_{2,j}-s_{1,j})$
where we have decomposed $b(\varphi)$ as a product of the generators $a_{i}$, $(b_{i}^{-1}a_{i}b_{i})^{-1}$, $\sigma_{j}$ and $z_{j}$, and $k_{gen}$ is the sum of the exponents of all the $a_{i}$ and $(b_{i}^{-1}a_{i}b_{i})^{-1}$ in this decomposition; $k_{\sigma}$ is the sum of the exponents of all the $\sigma_{j}$, and $k_{j}$ is the sum of the exponents of $z_{j}$ (according to the relations between the generators, $k_{gen}$, $k_{\sigma}$ and the $k_{j}$ do not depend on the decomposition of $b(\varphi)$ as a product of generators). Observe that this expression is homogeneous in $(v_{1},v_{2})$, i.e. $|\mathfrak{f}_{\kappa v_{1},\kappa v_{2}}(\varphi)|=\lvert\kappa\rvert\lvert\mathfrak{f}_{v_{1},v_{2}}(\varphi)\rvert$. Therefore, the maximum has to be attained when $s_{1}$ or $s_{2}$ is maximal, i.e. equal to $(k+1)A-1$. Since permuting $v_{1}$ and $v_{2}$ does not change $\lvert\mathfrak{f}_{v_{1},v_{2}}(\varphi)\rvert$, we can assume that $10\leq s_{1}\leq s_{2}=(k+1)A-$ 1. Let $k_{max}:=\max\{k_{j}\}$, $k_{min}:=\min\{k_{j}\}$, and let $j_{max}$ and $j_{min}$ be indices such that $k_{j_{max}}=k_{max}$, and $k_{j_{min}}=k_{min}$. Then,
$(k+g)\mathfrak{f}_{v_{1},v_{2}}(\varphi)\leq\frac{s_{1}-s_{2}}{2(k+2g-1)}(2k_{gen}-k_{\sigma})+\sum\limits_{s_{2,j}-s_{1,j}\geq 0}k_{max}(s_{2,j}-s_{1,j})+$
$+\sum\limits_{s_{2,j}-s_{1,j}<0}k_{min}(s_{2,j}-s_{1,j})\leq\frac{s_{1}-s_{2}}{2(k+2g-1)}(2k_{gen}-k_{\sigma})+s_{2}k_{max}-s_{1}k_{min}$
$=s_{2}\left(k_{max}-\frac{2k_{gen}-k_{\sigma}}{2(k+2g-1)}\right)+s_{1}\left(\frac{2k_{gen}-k_{\sigma}}{2(k+2g-1)}-k_{min}\right)$
with equality when $s_{1,j}=\delta_{j,j_{min}}s_{1}$ and $s_{2,j}=\delta_{j,j_{max}}s_{2}$.
Similarly,
$(k+g)\mathfrak{f}_{v_{1},v_{2}}(\varphi)\geq\frac{s_{1}-s_{2}}{2(k+2g-1)}(2k_{gen}-k_{\sigma})+\sum\limits_{s_{2,j}-s_{1,j}\geq 0}k_{min}(s_{2,j}-s_{1,j})+$
$+\sum\limits_{s_{2,j}-s_{1,j}<0}k_{max}(s_{2,j}-s_{1,j})\geq\frac{s_{1}-s_{2}}{2(k+2g-1)}(2k_{gen}-k_{\sigma})+s_{2}k_{min}-s_{1}k_{max}$
$=s_{2}\left(k_{min}-\frac{2k_{gen}-k_{\sigma}}{2(k+2g-1)}\right)+s_{1}\left(\frac{2k_{gen}-k_{\sigma}}{2(k+2g-1)}-k_{max}\right)$
with equality when $s_{1,j}=\delta_{j,j_{max}}s_{1}$ and $s_{2,j}=\delta_{j,j_{min}}s_{2}$. Since those expressions are linear in $s_{1}$, the extremal values are attained for $s_{1}=0$ or $s_{1}=s_{2}$. Set:
$R=k_{max}-k_{min},\,S=k_{max}-\frac{2k_{gen}-k_{\sigma}}{2(k+2g-1)},\,T=\frac{2k_{gen}-k_{\sigma}}{2(k+2g-1)}-k_{min}$
In the end, we get:
$\max\limits_{(v_{1},v_{2})\in V^{2}}\lvert\mathfrak{f}_{v_{1},v_{2}}(\varphi)\rvert=\frac{(k+1)A-1}{k+g}\max\left\{R,S,T\right\}$ (24)
and therefore
$\lVert\varphi\rVert\geq\frac{(k+1)A-1}{2(k+g)}\max\left\{R,S,T\right\}$
5 Hofer norms for braid groups on surfaces with boundary
Chen, in [chen23], proved non-degeneracy of the pseudonorms defined by the first author in [M]. We aim to show how Chen’s proof may be adapted to extend his theorem to:
Theorem 5.1.
Let $\underline{L}$ be a pre-monotone link on $\Sigma_{g,p}$, a compact symplectic surface of genus $g$ and with $p$ boundary components. Then there exists an $\varepsilon>0$, only depending on $\underline{L}$ , such that if $\varphi,\psi\in\text{Ham}_{\underline{L}}(\Sigma_{g,p})$ are such that $d_{H}(\varphi,\psi)<\varepsilon$, then $b(\varphi)=b(\psi)$.
This theorem in turn implies that the pseudonorms we defined are non-degenerate.
Corollary 5.2.
If $(\varphi_{i})\subset\text{Ham}_{\underline{L}}(\Sigma_{g,p})$ is a sequence of Hamiltonian diffeomorphisms preserving $\underline{L}$, and $\lVert\varphi_{i}\rVert\to 0$, then there exists a positive integer $n$ such that for every $i>n$ the braid $b(\varphi_{i})$ is trivial.
We are now going to adapt Chen’s proof to our setting.
In [chen23], the author considers Hofer-close Hamiltonian diffeomorphisms $\varphi$ and $\psi$. Choosing generating Hamiltonians for both of them and perturbing them gives rise to the associated Floer complexes, and continuation maps between them. If the generating Hamiltonians are $\mathcal{C}^{0}$-close, one may furthermore prove that the pseudo-holomorphic curves appearing in the continuation maps do not intersect the diagonal of the symmetric product, nor the divisor $D=z+\text{Sym}^{k-1}(\mathbb{S}^{2})$, where $z$ is the north pole of the sphere (outside the image of the embedding $\mathbb{D}\hookrightarrow\mathbb{S}^{2}$). The fact that the two Hamiltonians may be chosen to be close of course follows from the definition of the Hofer norm. One may then use these pseudo-holmorphic curves to produce a braid isotopy between chosen representatives of $b(\varphi)$ and $b(\psi)$ using the pseudo-holomorphic curves of the continuation maps.
In our case we wish to prevent pseudo-holomorphic curves from intersecting the diagonal and exiting the surface $\Sigma_{g,p}$: we glue discs of appropriate areas to the boundary components, so that $\underline{L}$ becomes monotone, and consider the centres $\zeta_{i}$, $i=1,\dots,p$ of the discs we just glued. Define the divisors $D_{i}=\zeta_{i}+\text{Sym}^{k+g_{1}}(\Sigma_{g})$: if $u:\mathbb{D}\rightarrow\text{Sym}^{k}(\Sigma_{g})$ is a pseudoholomorphic disc with Lagrangian boundary conditions on the link $\underline{L}$, if $[u]\cdot D_{i}=0$ for each $i$ then the image does not leave the image of the embedding $\Sigma_{g,p}\hookrightarrow\Sigma_{g}$, by positivity of the intersections between (pseudo)holomorphic submanifolds.
As in [chen23], we then decompose the differential in the Floer complex as a sum of contributions, each of them counting pseudo-holomorphic discs with fixed intersection number with the diagonal $\Delta$ and the divisors $D_{i}$. The function $\partial_{00}$, counting pseudo-holomorphic discs not intersecting $\Delta$ or any of the divisors $D_{i}$ is a differential, and we consider the complex $(CF(\underline{L},H),\partial_{00})$. As Chen points out, it is possible to do the same with continuation maps: if $K$ is a compactly supported Hamiltonian generating $\psi\in\text{Ham}_{\underline{L}}(\Sigma_{g,p})$, let $h:CF(\underline{L},H)\rightarrow CF(\underline{L},K)$ be the continuation map associated to a regular homotopy between $H$ and $K$. One may consider $h_{00}$, obtained from $h$ counting contribution of pseudo-holomorphic curves not intersecting $\Delta$ or any $D_{i}$: it turns out to be a chain homotopy between the two complexes $(CF(\underline{L},H),\partial_{00})$ and $(CF(\underline{L},K),\partial_{00})$. The arguments used in [chen23] to prove existence of pseudo-holomorphic curves contributing to $h_{00}$ (assuming that $\lVert H-K\rVert_{(1,\infty)}$ is small) carry over to our case: they in fact either involve local considerations around the Lagrangians, or use monotonicity of the link which we have now proved (see Proposition 2.1). As a consequence, one may follow through the proof of Theorem 1 in [chen23], which provides the braid isotopy we are looking for: such an isotopy is given by gluing a pseudo-holomorphic curve contributing to $h_{00}$ with a disc whose image does not leave the torus $\text{Sym}^{k+g}\underline{L}$. This completes the proof of Theorem 5.1.
6 A proof using Künneth formula
In this section we are going to present how, assuming the Hamiltonian diffeomorphisms at play are supported in a disc contained in the surface, one can prove Theorem 1.2 combining directly the results from [M] and [MT]. The former reference contains a statement about the Hofer geometry on the disc, and the latter one shows how to deduce results on the Hofer geometry of a surface with genus and boundary.
Assume a disc $D$ embedded in $\Sigma_{g,p}$, ($g,p\geq 1$) contains all the contractible components of a pre-monotone link $\underline{L}$, and does not intersect the non-contractible ones.
Denote by $\text{Ham}_{\underline{L},D}(\Sigma_{g,p})$ the set of Hamiltonian diffeomorphisms $\varphi$ satisfying:
  • $\varphi(\underline{L})=\underline{L}$;
  • the support of $\varphi$ is contained in $D$.
We have a map $\text{Ham}_{\underline{L},D}(\Sigma_{g,p})\rightarrow\mathcal{B}_{k}$ which is a surjective morphism of groups.
Theorem 6.1.
Let $\varphi\in\text{Ham}_{\underline{L},D}(\Sigma_{g,p})$; denote by $h$ its braid type. Then,
$\lVert\varphi\rVert\geq\frac{1}{4(k+g)}\frac{(k+1)A-1}{(k+2g-1)}\lvert\mathrm{lk}(h)\rvert$
Proof.
For $s\geq 0$, if $\Sigma_{g}$ has non-empty boundary, we choose a symplectic embedding of $(\Sigma_{g,p},\omega)$ into a closed surface $(\Sigma_{g}(s),\omega^{s})$ of genus $g$ and area $1+s$, so that $\underline{L}$ becomes an $\eta_{s}$-monotone link, where
$\eta_{s}=\frac{(k+1)A-(1+s)}{2(k+2g-1)}$
We get a Hofer-Lipschitz spectral invariant $c^{s}_{\underline{L}}$, defined for a Hamiltonian $H:S^{1}\times\Sigma_{g}(s)\to\mathbb{R}$ by
$c^{s}_{\underline{L}}(H)=\frac{1}{k+g}c^{s}_{\text{Sym}\underline{L}}(\text{Sym}^{k+g}H)$
where $c^{s}_{\text{Sym}\underline{L}}$ is the Lagrangian spectral invariant of $\text{Sym}\underline{L}\subset\text{Sym}^{k+g}(\Sigma_{g}(s))$. We write $\underline{L}=L_{1}\cup...\cup L_{k}\cup\alpha_{1}\cup...\cup\alpha_{g}$, where the $\alpha_{i}$’s are the non-contractible components, and define $\underline{K}:=L_{1}\cup...\cup L_{k}\subset D$. One can view $\Sigma_{g}(s)$ as a connected sum $\mathbb{S}^{2}\#E_{1}\#...\#E_{g}$, where the sphere $\mathbb{S}^{2}$ contains $D$ (and therefore $\underline{K}$), and the $E_{i}$ are copies of the 2-dimensional torus, with $\alpha_{i}\subset E_{i}$. Now, suppose $H$ is supported in $D$, and denote by $\underline{L}^{\prime}$ (resp. $\underline{K}^{\prime}$) the image of $\underline{L}$ (resp. $\underline{K}$) by $\phi^{1}_{H}$. By taking a small perturbation of $H$ sending the $\alpha_{i}$ to some transverse $\alpha_{i}^{\prime}$, and applying $g$ times the Künneth formula [MT, Theorem 6], we get an isomorphism of filtered chain complexes
$CF^{*}(\text{Sym}\underline{L},\text{Sym}\underline{L}^{\prime})\cong CF^{*}(\text{Sym}\underline{K},\text{Sym}\underline{K}^{\prime})\otimes\bigotimes\limits_{1\leq i\leq g}CF^{*}(\alpha_{i},\alpha_{i}^{\prime})$
On the right hand side, $\underline{K}$ and $\underline{K}^{\prime}$ are links in $(\mathbb{S}^{2},\omega^{s}_{\mathbb{S}^{2}})$, and the $\alpha_{i},\alpha_{i}^{\prime}$ are circles in $(E_{i},\omega^{s}_{i})$, where the symplectic forms satisfy the following:
  • $\omega^{s}_{\mathbb{S}^{2}}$ coincides with $\omega^{s}$ on $D$ (which is away from the connected sum region), and has total area $1+s+4g\eta_{s}$;
  • $\omega^{s}_{i}$ coincides with $\omega^{s}$ on a neighbourhood of the homotopy between $\alpha_{i}$ and $\alpha_{i}^{\prime}$.
Hence, we get that $c^{s}_{\text{Sym}\underline{L}}(\text{Sym}^{k+g}H)=c^{s}_{\text{Sym}\underline{K}}(\text{Sym}^{k}H)$, and therefore
$c^{s}_{\underline{L}}(H)=\frac{k}{k+g}c^{s}_{\underline{K}}(H)$
where the right hand side is computed in $(\mathbb{S}^{2},\omega^{s}_{\mathbb{S}^{2}})$. Since this holds for any $H$ supported in $D$, it holds for a Hamiltonian generating $\varphi$ and for its iterates, so we get after homogenisation (and renormalisation)
[…]
(it is shown in [CGHMSS22, Section 7.3] that the (renormalised) homogenised spectral invariants do not depend on the choice of $H$ generating $\varphi$). By the Hofer-Lipschitz property of spectral invariants, we obtain
$2\lVert\varphi\rVert\geq\lvert\mu^{0}_{\underline{L}}(\varphi)-\mu^{(k+1)A-1}_{\underline{L}}(\varphi)\rvert=\frac{k}{k+g}\left(\lvert\mu^{0}_{\underline{K}}(\varphi)-\mu^{(k+1)A-1}_{\underline{K}}(\varphi)\rvert\right)$
Now, according to [M, Theorem 1.4],
$\lvert\mu^{0}_{\underline{K}}(\varphi)-\mu^{(k+1)A-1}_{\underline{K}}(\varphi)\rvert=\frac{1}{k}\eta_{0}\lvert\mathrm{lk}(h)\rvert=\frac{1}{2k}\frac{(k+1)A-1}{(k+2g-1)}\lvert\mathrm{lk}(h)\rvert$
and therefore we get the theorem. ∎
[…]